A Complete Bibliography of Publications in the Collnet Journal of Scientometrics and Information Management

Nelson H. F. Beebe University of Utah Department of Mathematics, 110 LCB 155 S 1400 E RM 233 Salt Lake City, UT 84112-0090 USA

Tel: +1 801 581 5254 FAX: +1 801 581 4148

       E-mail: beebe@math.utah.edu, beebe@acm.org,
                  beebe@computer.org (Internet)
      WWW URL: http://www.math.utah.edu/~beebe/

10 October 2024 Version 1.12

Title word cross-reference

3 [HH20]. ^{nd} [Asu11]. ^{st} [Ano22a, Glä11, Mar20]. ^{th} [Ano18, Ano19a, Ano19b, Ano22a, Asu11, DHZK09, MK15, MKP19, Mar20, Mus07, YLM20]. h [Ano22k, MR16a, LM21, Rah20, RS18, RL08, kWhHR07, YBL17].

-based [YBL17]. -index [LM21, Rah20, RS18, RL08, kWhHR07]. -indices [MR16a]. -Kinase [HH20]. -type [Ano22k, RL08].

//www.slp.org.in/collnet2015 [KJ16].

1 [KK08]. 11th [KJ16]. 12th [KLJ17]. 14 [Ano18, MKP19]. 15 [Ano19a, YLM20]. 16 [Ano19b, Ano22a, Mar20]. 16th [KJ16]. 17th [KLJ17]. 19 [Ano22q]. 1989-2019 [JSM20]. 19th [MKP19].

2 [KdBKK08]. 2.0 [Sin15]. 21 [Glä11]. 25010 [DSM20]. 26-28 [KJ16].

5 [DHZK09].

6 [Asu11].

Abstracts [Ver09]. academia [MES15]. Academic [HM10, LC15, cSL14, SKNN08, Ano22e, CWS20]. Academy [CF13, MLJM14]. acceptance [Ano22x]. Access [RC13, Ano22b, Ano22p, EE20, GR21]. account [kWhHR07]. Accounting [HAM^{+}21]. Active [YLZS15]. Activities [Cho12, ES15, SKNN08]. Activity [BP10, KhJ13, MLMJ13]. address [RVB^{+}14]. adhesive [KZZV20]. Administering [Pal16]. adoption [Ano22x, MJC19]. advancement [MG18]. affect [BH12]. affected [ZKJK19]. Africa [EE20]. African [AE21, LOA16]. against [dJSPR21]. Agencies [AO07]. agreement [ÖP12]. Agricultural [MG12, RS11, GR21]. Agriculture [MG12, UY17]. Aguillo [Agu08a, Agu08b]. Ahead [BS12]. Ahvaz [AOR14]. Ahvaz-Iran [AOR14]. AI [Ano22s]. AIDS [GBK11]. AIDS/HIV [GBK11]. Alexa [Vau12]. Algorithms [ES16b]. Alternative [Fak14, Vau12, Ano22q, ME12]. altmetric [SR17]. Ambedkar [SK11]. among [AAMH17, Ano22r, AO07, hJjJiK13, KGB11, LLO09, MS12a, Rao12, Sci14]. Analysis [CKM16, CWLL11, ES16a, EE20, dCPFA10, GDY14, GRAY15, GBK11, GB11, GBG13, GG14, hJjJiK13, JKK13, Kal16, KGBN12, KG09, KMG11, KK07b, KK17b, KhJ13, iKhJ14, LGJ14, LC15, MR16b, OKRS14, PK15, RC13, RR15b, RVB^{+}14, SS14, SA07, SK11, Vau12, YLZS15, Ahm18, Ano22f, Ano22h, Ano22l, Ano22t, Ano22x, HB21, CKJC13, Dar12, DK17, EH18, HH20, HAM^{+}21, HYWH20, JWH20, JLH21, NK20, LPGC20, LFH18, MG18, MS07a, MS12b, MHS16, MPPH17, BM21, MIS^{+}20, Naz18, Neg21, OB10, ÖP12, PRP21, RP21, RRB17, RS18, RSM^{+}19, GR21, SMM08, SAPM13, SA16, dJSdJP21, Sha19, SM18, SB14, ES20, LS19, YS20, UE19, WWT21, WH19, HW20, WWH20, GW21, CZ20]. analytical [AE21]. Analyzing [ES15, ES16b]. Announcement [Ano18, Ano19a, Ano19b, Ano22a]. Annual [DM14]. APC [Ano22b]. apoptotic [JLH21]. Applicability [Fak14, RVF^{+}13, YLX14]. Applicable [KK07b]. Applicants [hJjJiK13]. applicateness [RVF^{+}13]. Application [BVNS21, MJC19]. Applications [MR16a, PRP21, SNK11, Ano22n, BK17, MG18]. Applied [UE19, LBN21]. applying [Lev07]. Approach [MT08, Dar12, HAN14, RVF^{+}13, SJMN16]. Approximation [PB18]. Archaeology [Hér08]. Area [Tij08, Sud10]. Art [Hér08]. Article [PPB09, Ano22b]. Articles [dCPFA10, PG09, Ano22n, EH18, JLH21, Lev07, MFH18]. Artificial [LBN21, MRSE21, HW20]. ASE [ME12]. Asia [GR21]. Asian [SM18, EJHG13, JBM10, SA07]. ASME [Ano22c]. aspects [Egg13, Mus17]. Assessing [BP10, Dar12, Rah20]. Assessment [AAFH14, NK14, AWAMA18, Ano22e, Ano22g, Ban19, BK17, DPR^{+}18, DGG18, GD19a, GD19b, VHA21]. Assignees [CF13]. associated [dJSdJP21]. Association [SM18, DSM20, JAAN14]. Associations [LLO09]. assurance [DSM20]. Attention [GRAY15, Rou12]. Australian [BWC08]. Author [GKL08, KK07a, PKL09, PG09, SJM16, Ano22m, Ano22k, KK08, OM08, TKSB18, WKK10, kWhHR07, KdBKK08]. Authored [Mus14]. Authors [ES16b, RS08, Rao12, Sci14, MR19, WWT21, kWhHR07]. Authorship [SDK14, VR15, Ano22k, GK19, Glä11, KK08, KdBKK08, MR18, OM11, OKRS14, OCK17]. Automated [MT08]. Automatic [LR14].

Babasaheb [SK11]. Baidu [GRAY15]. Barriers [OAR14, Saf14]. Based [GRAY15, MR16a, Ano22d, Ano22q, Ano22r, Ano22k, BP10, CKJC13, DM14, Dis15, DSM20, FPK21, HS15, MW12, MJC19, LM21, MP17, MR16b, OKRS14, RS18, RVF^{+}13, MS21, Sud10, YS19, YS20, SR17, YBL17, CZ20, MHS16]. basis [VB17]. Battery [hJjJiK13]. beauty [KK17a]. behavior [BH12, Zit13]. behavioral [MRAA20]. Bengal [RS11]. between [AT15, Ano22h, Ano22r, HA15, iKhJ14, NKK15, RVB^{+}14, SCdM^{+}12, Zit13]. Biblio [MS12a]. Biblio-Blogosphere [MS12a]. Bibliographic [Ano22c]. Bibliometric [Ano22d, AM10, ACO21, CWLL11, HAM^{+}21, HYWH20, KGB11, KK07b, MLJM14, MS14, Mus17, MIS^{+}20, MR16b, NL16, PPB09, Ram14, RT15, RS11, STP18, Sin15, ES20, Sun07, YS19, YS20, Ver09, VT19, GW21, Ano22f, Ano22j, Ano22l, Ano22q, Ano22x, Ban19, GB21, DPR^{+}18, DKK18, EH18, EE20, HH20, Ho18, JWH20, JLH21, LFH18, LBN21, MS07a, MS12b, MRAA20, MPPH17, BM21, Naz18, PRP21, RP21, dJSdJP21, MS21, Sha19, Tij08, UE19, WH19, HW20, WWH20]. Bibliometrics [MG20, Glä08, LGJ14, MFH18, Ela19, GK19]. bibliometrics/scientometrics [GK19]. bilingual [Hol11]. bioaccumulation [LFH18]. biochemistry [NL16]. Biographical [Agu08a, Cha07a, Hus09, Mey07a, Osa08, Öze11, San10, Sha10, Zey09]. Biological [KK07a]. Biotechnology [BP10]. Bloggers [YLZS15]. Blogosphere [MS12a]. Blogs [YLZS15]. Board [LC15]. Bradford [BVNS21, Sud10]. Bradfordizing [May08]. Brazil [dCPFA10, SCdM^{+}12, dSVS12]. Brazilian [Ano22d]. Brief [Glä11]. Bulgaria [VHA21]. Buonocore [KZZV20]. Buruli [Ban19]. Business [HAM^{+}21].

CAGR [Ela19]. Calculating [RR14, kWhHR07]. Calculation [Ela19]. Call [Ano18, Ano19a, Ano19b, Ano22a, KN11]. Canada [MES15]. Canadian [ES15]. cancer [MRSE21, BM21, MP17, OCK17]. capabilitie [MPPH17]. Car [KhJ13, iKhJ14]. carpentry [Ano22j]. CAS [CF13]. Case [BP10, CF13, KS10, LG09, Mus14, RM09, RS08, RVB^{+}14, Ano22s, Bur09, Ela19, FPK21, GK16, MS13, MH17, Mus07, Saf14, MS21, ZKZ19]. case-study [MS13]. Categories [ME12, Ano22r]. category [YS19, WH19]. Cell [iKhJ14, KGBN12]. cells [dJSdJP21]. Center [Cho12]. Central [HM10]. centralities [Ano22c]. Centrality [KK07b]. centre [Pal16]. Century [Gar12, Glä11]. Chagas [dJSdJP21]. Chamran [AOR14]. Chandigarh [KMG11, LG09]. Chandra [Cha07a, Cha07b]. change [SS19]. Characteristics [SB14]. Characterization [cMhLyW14, OCK17, SB14]. charges [Ano22b]. charting [CSN20]. Chemical [SMM08, SM09]. Chemistry [KK07a, OM08, ES20]. China [BS12, LMvI08, Mus07]. Chinese [CF13, HLK^{+}09, LS19, ZKJK19]. chip [BM21]. Citation [Bur09, CKJC13, Gar12, HM10, Kal16, KK17b, MHS16, OKRS14, PK15, PG09, RC13, RVB^{+}14, SK11, SWH19, YLC13, ASW^{+}19, Ano22m, EBS16, Ho18, NK20, LOA16, OM08, PB18, SB14, SR17, YBL17, Zit13]. Citations [DM14, Rao12, Sci14, ZS12, Bar15, LM21, Pea20]. Cited [BBS18, PPB09, SGP16, EH18, HH20, HYWH20, MFH18, Naz18]. Citedness [GP15]. CiteScore [OAdS21]. CiteSpace [CZ20]. citing [Zit13]. citing-side [Zit13]. Clark [RY18]. Class [AA09]. Classic [Ho18, JLH21, SWH19]. Classification [LR14]. Climate [SS19]. cloud [ACO21, Bha17, EE20]. clusters [OAdS21]. Co [Glä11, GKL08, KK07a, KK08, KdBKK08, OKRS14, PKL09, WKK10, Ano22m, GK19, OB10, OM08, OCK17, ÖP12, RSM^{+}19, kWhHR07, KdBKK08]. Co-Author [GKL08, KK07a, PKL09, KK08, WKK10, Ano22m, KdBKK08]. co-authors [kWhHR07]. Co-authorship [Glä11, KK08, KdBKK08, OKRS14, GK19, OCK17]. co-citation [OM08]. co-occurrence [RSM^{+}19]. co-publication [OB10]. co-word [ÖP12]. cognitive [ZKJK19]. Coherence [RM09]. Coldlib [MT09]. collaborating [Bar15]. Collaboration [AST^{+}12, Bas14, dB13, CF13, EGK08, ES16b, EJHG13, HLK^{+}09, HWL^{+}13, KK12, KKS14, LLO09, RN17, SM09, dSVS12, AOR14, AO07, GK19, KK17a, MS07b, MES15, MR18, Mus07, Naz18, OB10, RSM^{+}19, SA16]. Collaborations [Ron11]. Collaborative [Sun07, VR15, SJM16]. College [KMG11, MS14]. COLLNET [Ano18, Ano19a, Ano19b, Ano22a, Asu11, DHZK09, KN11, KJ16, KLJ17, MKP19, Mar20, YLM20, AMN17, KK12, MT09, GB21, RR15b]. collnet2015 [KJ16]. Columbia [SDK14]. combinatoric [RSM^{+}19]. Commemorating [KK12]. commerce [Ano22h]. commercialization [LS19]. Communication [Glä11, Har14, Hér08, NK20, LR14, MS12a, RC13]. Communications [RRA14]. Community [MT09]. Comparative [SGP16, HAN14, KGBN12, NK20]. comparing [Ela20]. Comparison [AAMH17, Bas14, MS21, YLC13, LM21, MES15]. Competency [BP10]. Competitive [SNK11, MLA12, MLMJ13]. Competitiveness [hJjJiK13]. components [MH17]. compounds [MFH18]. Computer [ES16b, OM11, ES16a, STP18]. Computers [GRAY15]. computing [ACO21, Bha17, EE20, GD19b]. Concept [Ali10, MT09]. Conference [Ano18, Ano19a, Ano19b, Ano22a, Asu11, DHZK09, KN11, KJ16, KLJ17, MK15, MKP19, Mar20, YLM20]. Conferences [ES16b, AMN17, ES16a]. Conjunctivitis [GBG13]. Consensus [AAFH14]. Consistent [DM14].

Consolidation [JKK13]. Constructing [PPB09]. content [DV18]. context [Zit13]. Contribution [Ahm18, GK19, HA15, KG09, LG09, cSL14, MR19]. Contributions [RT15, SM09, LOA16, SK17]. Cooperation [CWLL11, Mus11, SM18, OAR14]. core [ASW^{+}19]. Corona [Ano22q]. corpora [MS13]. corpus [MSH08, MS12b]. correlate [SR17]. Correlates [GP15]. correlation [Ano22o]. Corruption [RP21]. count [Ano22m]. Counting [SNK11]. Countries [EJHG13, JBM10, KGB11, RRA14, SA07, SM18, BMNS13, Bar15, BCB13, AE21, EKM^{+}18, KGBN12, MS07b]. Country [Bas14, BP10, Mus14, RS18, BBS18, Kok19, NKK15]. coupling [Ano22c]. Covid [Ano22q]. Covid-19 [Ano22q]. credit [Rah20]. critical [Ano22e]. Cross [SA07]. Cross-National [SA07]. cryptology [Pal16]. crystals [KK17a]. Cultures [Hér12].

dance [WH19]. Data [CKM16, KK07b, Vau08, Vau12, Ano22d, Ano22j, EE20]. Database [BMNS13, CKJC13, DPR^{+}18, MR16b, RSZ15, Dis15, MP17, MIS^{+}20, Shu19, VT19]. databases [Lev07]. DB [MS07a]. DBs [MS12b]. Debrecen [DSM20]. Decade [KK12]. decades [JSM20]. December [KLJ17, MKP19]. Defence [MR16b]. Definitions [MR16a]. degradation [MG20]. Delhi [KJ16, MS14]. Dengue [DKG10]. dentistry [KZZV20]. departments [FPK21]. DESIDOC [VR15]. design [DV18]. Desk [Ano22i, Hus07a, Hus07b, HK08a, HK08b, HK09a, HK09b, Rao13, HK10, HRK10, Kre19, RK11a, RK11b, RKÖG12, RK12]. detection [ACO21]. determinants [Kok19]. Developing [BP10, Lev07]. Development [AT15, HA15, cSL14, RSM^{+}19, Ano22f, Ano22q]. Developments [Har14, RL08]. devices [BM21]. diagnosis [BM21]. dialysis [HB21]. Diaspora [Bas14]. diasporas [Sha19]. Differ [ZS12]. differences [Ano22j]. different [Ano22d, BWC08]. diffusion [Dar12]. Digital [Har14, KS10, DSM20, GR21]. directions [CSN20]. Disciplinary [Hér12, RM09]. discipline [CSN20]. disciplines [RVB^{+}14]. Discovering [MRSE21]. discovery [Ano22s]. Discrepancies [ZKZ19]. disease [dJSdJP21]. disorder [Shu19]. Dispersion [Hér12]. Dissertations [Hér08, UY17]. Distance [PPB09]. Distinct [OAdS21]. Distribution [GKL08, KK07a, KK08, KdBKK08, PKL09, SDK14, OM11, PB18]. Distributions [KK08, RS08, PB18]. Diversity [KhJ13, RM09]. Division [KS10]. Do [SR17, ZS12]. Doctoral [Sud10]. Documents [YLX14]. Does [Mus11]. domain [Ano22u, SJMN16]. Domains [MLJM14]. Dots [Ano22n]. downloads [Ano22r]. Dr. [Agu08a, Agu08b, Cha07a, Cha07b, KKYR17, Öze11, SK11]. Drawbacks [Rah20]. DSRS [SJMN16]. During [EJHG13, Ano22v, Ban19, DGG18, ES16a, EKM^{+}18, GBK11, GB11, GBG13, GG14, GD19a, GD19b, JSM20, JMS19, KG09, LG09, MJC19, MR16b, NL16, OM11, OCK17, SP20, MR19, RSZ15, YS20]. dye [Ano22w]. dynamic [MRSE21, MPPH17]. Dynamical [Egg13]. Dynamics [HWL^{+}13]. e-commerce [Ano22h]. e-Government [Ano22f]. E-Journal [JKK13]. E-learning [MJC19, Neg21]. Early [BM21]. East [BCB13, EKM^{+}18]. Eastern [NL16, BMNS13]. EBL [MHS16]. Economic [AT15, KJ16, LS19]. economics [FPK21, JSM20]. Editor [Ano22i, Hus07a, Hus07b, Kre19, RK11a, RK11b, RKÖG12, RK12, Rao13]. Editorial [LC15]. Editors [HK08a, HK08b, HK09a, HK09b, HK10, HRK10]. Education [Ali10, Ali11, MLJM14, Ano22t]. Effects [May08, WX14]. eigenfactor [AAMH17]. Eight [SGP16]. Eighth [KN11]. Electronic [GD19a]. Elman [PPB09]. Embrapa [dCPFA10]. Emerging [Muk11]. eminence [WZ17]. empirical [DK17]. Endocrinology [EKM^{+}18]. Energy [iKhJ14, PKL09]. Engine [MT08]. Engineering [RR15a, SGP16, SG10, AAMH17, Kok19, ZKZ19]. Engines [ME12, MG12]. English [ZS12]. environment [MS07a]. Environmental [Muk11, AAMH17]. era [KS10]. Erdos [Glä11]. errors [Ela19]. ESI [YS19]. ESL [Pea20]. estimating [Lev07]. Estimation [SJMN16]. Estonia [NK14]. EU [MS07b, Mus07, Mus17]. Eugene [KKYR17, Rou12, SS10]. Europe [LMvI08]. European [Tij08]. EuroSys [ES16a]. Evaluating [ASW^{+}19, PG09]. Evaluation [KS10, MK12, Muk11, Nam17, Ron11, SNK11, DV18, JSM20, May08, MS21, VHA21]. Evidence [OP10, VT19, ZKJK19, MHS16]. Evidence-based [MHS16]. Evidences [BS12]. Evolution [CF13, MRSE21]. Examination [Ano22h, cSL14]. Examining [ÖP12]. expanded [Ho18, SWH19, HM10]. Expected [BH17]. Experience [MG12]. Experiences [HA15]. Experimental [KKS14]. Explained [KK07a]. exploration [CZ20]. Explorations [Hér12]. Exploratory [YLC13]. Explore [YLX14]. Exploring [FPK21]. express [Sha19]. expressions [ÖP12]. extensive [DSM20]. External [SNK11].

F [Agu08a]. Faces [dB13]. facial [ÖP12]. Factor [MMM09, PPB09, Adi19, AAMH17, AA17, MSH08, Rah20, VB17, Fak14, GP15]. Factors [BWC08, BH12, Sci14, AOR14]. facts [Glä08]. faculty [AOR14, LM21, Saf14]. farming [Ano22v]. Federal [MLMJ13]. feedback [Pea20]. festschrift [Rou12]. Festschrifts [Rou12]. fiction [Glä08]. Field [RT15, ACO21, Bha17, DK17, EH18, EKM^{+}18, JAAN14, MS07b, MFH18, Neg21, STP18, SA16, GW21]. Fields [hJjJiK13, KhJ13, Muk11, BTML16, Zit13]. figures [ÖP12]. Film [KS10]. Films [KS10]. financial [CSN20]. firms [ZKJK19]. first [Ano22k]. flow [ES20]. flows [AG17]. focused [BSdMAM15]. Food [CKM16, PRP21]. forecasting [SJMN16]. Foreign [MR18]. forensic [RRB17]. Foresight [PNA15]. Formalizing [CKM16]. Formation [KKS14]. FP [Mus17]. Fractional [SNK11]. Framework [Mus07, HAN14]. France [KLJ17]. French [Hér08]. Frequencies [GKL08, KK07a, PKL09, KdBKK08, WKK10]. Front [Ano07a, Ano07b]. Frontier [Vau08]. Fuel [iKhJ14]. Funded [Kok19]. Funding [Ron11, SNK11, MLA12, MLMJ13]. future [Ano22n, CSN20].

G15 [KGB11]. G8 [RRA14]. Garfield [KKYR17, Rou12, SS10]. Gauging [MRAA20]. Gender [Ano22j, KdBKK08]. General [Hér08]. Generation [MT08]. genetic [Shu19]. Genetics [SAPM13, SMA15, SA16]. geography [SB14]. geology [JMS19, RS18]. Gestalt [KK07a]. GH} [Ano22g]. GH}-index [Ano22g]. Global [AST^{+}12, BS12, DKG10, GD19a, JWH20, MT09, AG17, Ano22l, Ano22t, Ano22u, Ano22v, Ano22w, Ban19, BJ15, Bha17, DGG18, MG20, LFH18, SMA15, SS19]. GMCH [KMG11]. Google [LOA16, ZS12]. Government [Ano22f, KMG11]. Gramscian [Ano22l]. grant [BK17, RN17]. graph [MG18]. graphics [STP18]. great [KKYR17]. Greece [FPK21]. Green [KhJ13, iKhJ14, YS19]. groundnut [DKK18]. Group [CWLL11, WWT21, OAR14]. Growth [KKS14, KJ16, RSZ15, SMM08, SA16, SP20, LS19].

h5 [AAMH17]. h5-index [AAMH17]. Health [KA08, Ano22u, HYWH20]. hearing [RN17]. Hegemony [Ano22l]. Helicobacter [SWH19]. HHV8 [WWH20]. High [BP10, SSKM21]. Higher [Ali10, Ali11, MLJM14, Ano22t]. Highly [BBS18, Naz18, HH20, HYWH20, MFH18, SGP16]. Hill [NL16]. Hindawi [Ano22b]. Historical [Hér08, Hér12, SWH19]. History [Mus11]. HIV [GBK11]. holistic [HNN09]. Homage [KKYR17]. Hospital [KMG11]. Hosted [KJ16]. http [KJ16]. Humanities [JAAN14, AG17, Ano22d, TKSB18]. humble [KKYR17]. Hungary [Ano22e]. Husain [Hus09]. Hybrid [RM09]. Hyperlink [Vau12].

I3 [YBL17]. I3-type [YBL17]. Identification [KMA20]. Identifying [BTML16, CF13, ZKZ19]. IEC [DSM20]. IITs [DV18]. images [ÖP12]. Impact [AM10, BWC08, ES16b, EJHG13, Fak14, GP15, KTCA19, LMvI08, MLA12, MMM09, Mus14, PG11, RC13, Sci14, VHA21, Adi19, AAMH17, AWAMA18, AA17, Ano22m, EE20, FPK21, MLMJ13, MSH08, Rah20, VB17]. Impacts [Har14]. Implications [Mus11]. important [BTML16]. Increase [Bar15]. Index [AAFH14, Ano11, Ano12, DM14, Ela20, GRAY15, HM10, OKRS14, RR14, AAMH17, Ano22g, Ho18, LM21, Rah20, RS18, RL08, kWhHR07, SWH19]. Indexes [PG09, YLC13]. Indexing [Gar12, Ano22b]. India [KJ16, Bar15, DK17, GBK11, GB11, GD19b, JSM20, KGBN12, KA08, LMvI08, LM21, Mus07, RR15a, VB17, WKK10, WDG07]. Indian [AWAMA18, Ahm18, Bas14, HB21, BP10, DV18, DKK18, GK16, GK19, GG14, GD19b, HM10, JMS19, KG09, KTCA19, MW12, MR18, Muk11, MP17, PG11, PK15, RT15, MR19, SS14, SMM08, SM09, SM14, Shu19, SG10, Ver09]. indicator [AAMH17, Ela20, SNK11]. Indicators [AM10, HA15, ME12, RR14, SG10, dSVS12, ASW^{+}19, Ano22e, BH17, HAN14, MRAA20, Pal16, WZ17, YBL17]. Indices [Fak14, Ano22l, Ano22k, MR16a, RL08]. Individual [PG09, Ano22e]. Indonesia [DPR^{+}18]. Indonesian [Ano22f]. Industrial [cSL14]. Industry [WX14, MES15, MH17]. influence [SJMN16]. Influential [YLZS15, BBS18]. informatics [Ano22u]. Information [Ali10, Ali11, AA09, Cho12, Egg13, GKL08, HS15, Hér08, KK17b, NK20, LLO09, LC15, Mar20, MMM09, RR15b, SM18, UE19, VR15, ASW^{+}19, AG17, Ano22p, Ano22t, GB21, BVNS21, CKJC13, DK17, KKYR17, MW12, LM21, Rou12, Saf14, UY17]. Informetric [PG11, Vau08, RVF^{+}13]. Informetrics [Ano18, Ano19a, Ano19b, Ano22a, Asu11, Cho12, DHZK09, KN11, KJ16, KLJ17, MK15, MKP19, Mar20, YLM20]. Ingredients [CKM16]. inhibiting [AOR14]. innovation [Nam17]. insomnia [JWH20]. Institute [DV18, KJ16, MMM09]. Institutes [VB17]. institutional [BBS18]. Institutions [Bas14, BSdMAM15]. integrated [Tij08]. Integrating [HAN14]. Intellectual [AT15]. intelligence [MRSE21, HW20, LBN21]. Interactions [RRA14]. Interdisciplinary [cMhLyW14, OB10, Ron11, LPGC20]. International [Ano18, Ano19a, Ano19b, Ano22a, Asu11, Bas14, DHZK09, ES16a, HLK^{+}09, KN11, KJ16, KLJ17, MK15, MKP19, Mar20, Mus07, YLM20, LBN21, MR19, DHZK09]. Internationalizing [SKNN08]. internet [EE20, Vau12]. Inventions [WX14]. Investigating [OB10]. Investigation [RSZ15, ZS12, BBS18, EGK08]. investigations [HAN14]. IPC [iKhJ14]. Iran [AOR14, HNN09, HA15, MK12, Nam17, Neg21, OAR14, RRA14]. Iranian [AO07, AMN17, EGJ13, JAAN14, MG12, OM08, OKRS14, OCK17, RSZ15, ZS12]. ISI [MMM09, ZS12]. Islamic [HAM^{+}21, DPR^{+}18]. ISO [DSM20]. ISO/IEC [DSM20].

Japan [LMvI08, SKNN08]. Japanese [Sun07]. JASIST [RS08]. JCR [MS12b]. Journal [GB21, GP15, JKK13, KK07a, MSH08, PG09, RR15b, RS18, RVB^{+}14, SDK14, VR15, AAMH17, ASW^{+}19, AA17, Ano22r, EGK08, LLZ16, LBN21, MR19, SJMN16, STP18, SR17, Fak14, GP15, GKL08, HAM^{+}21, KKS14, MR16b, UE19]. journal-based [Ano22r]. Journals [KKS14, LC15, PK15, RC13, RS08, RSZ15, ZS12, AAMH17, AWAMA18, ASW^{+}19, Ano22c, Ano22b, Ano22p, Ano22r, NK20, KdBKK08, MW12, MS12b, MS13, MRAA20, MR18, OAdS21, RS18]. Jundishapur [AOR14].

KESLI [JKK13]. keywords [TKSB18]. Kientometrics [HAN14]. Kinase [HH20]. KISTI [Cho12]. Knowledge [CKM16, CSN20, Cho12, DHZK09, HAN14, MT09, MG12, MH17, RVB^{+}14, WX14, BH12, KMA20, Saf14, ZKJK19, Rou12]. Korea [CKJC13, Gup10, YLC13]. Korean [CKJC13, ÖP12]. KSCD [CKJC13]. KSHV [WWH20]. KSHV/HHV8 [WWH20].

L [San10]. lab [BM21]. lab-on-a-chip [BM21]. Laboratory [WKK10]. land [MG20]. landscape [MPPH17]. language [RN17]. Lanka [Dis15]. last [HB21, JSM20]. Law [SDK14, BVNS21, OM11, SDK14, YLX14, Sud10]. Laws [SS07]. Learning [AA09, Ano22p, Ano22x, BJ15, MJC19, Neg21]. Leimkuhler [BVNS21]. Leptospirosis [RT15]. level [BBS18, Bur09, LS19, Tij08, WWT21]. lexical [MS07a]. Librarian [MG12]. Librarianship [MHS16]. librarie [DV18]. Libraries [GD19a, SS07, Sin15, Ano22s]. Library [HS15, KS10, LC15, MK12, SM18, VR15, ASW^{+}19, Ano22p, DK17, DSM20, MW12, LM21, KK17b, NK20, LLO09, UE19]. life [CWS20]. lightweight [SJMN16]. linked [Ano22b]. Linking [Hol11, Vau12]. LIS [KK17b, MS12a, MR18, PK15]. LISAS [LLO09]. literacy [Ano22t, CSN20]. Literature [KK17b, SAPM13, SM14, Sin15, Ano22x, Bha17, BVNS21, Kok19, NL16, RRB17, SMM08, SMA15, SB14, UE19]. Literatures [Ram14]. Liu [KK12, CWS20]. Local [AST^{+}12]. Lotka [OM11, SDK14, YLX14]. Lymphoma [GG14].

M [Sha10]. Macau [MKP19, MKP19]. machine [Ano22p]. macro [Tij08]. macro-level [Tij08]. Main [MLJM14]. maize [YS20]. Major [ES16b, hJjJiK13]. Management [Ali10, Ali11, AA09, GDY14, FPK21, KMA20, MH17, GB21, RR15b]. Many [dB13]. Map [iKhJ14]. Mapping [Bha17, GBK11, GB11, MPPH17, OP10, RR15a, RS11, SM14, SG10, dSVS12, Ano22h, Ano22u, Ano22v, ACO21, BSdMAM15, CSN20, Neg21, OCK17, STP18, dJSdJP21, MS21, YS19]. Maps [MT08]. Marathwada [SK11]. Market [GRAY15]. Martin [Mey07a]. Materials [WDG07]. mathematical [MS12b, MS13]. Mathematics [Ano22j]. matrix [RSM^{+}19]. Matter [Ano07a, Ano07b, Mus11]. means [Pal16]. Measure [KK07b]. measures [Ano22d]. Measuring [Ano22m, SP20]. Media [Har14, NK20]. Median [GP15]. Medical [AOR14, KMG11, MS14, RC13]. Medicinal [Dis15]. medicine [Ano22q, KKS14]. medicine-based [Ano22q]. Meeting [Asu11, KJ16, KLJ17, MKP19, Ano18, Ano19a, Ano19b, Ano22a, DHZK09, KN11, Mar20, YLM20]. Members [LC15, AOR14, Saf14]. Memories [MS07a]. Memory [MS07a]. meta [RY18]. meta-table [RY18]. Metabolism [EKM^{+}18]. Method [Sha19]. Methodological [Iva08, Mus17]. methods [AO07, LLO09]. metric [Lev07]. Metrics [GRAY15, Ano22b, Ano22r]. Meyer [Mey07a, Mey07b]. Micro [Bur09]. microfluidic [PRP21]. Middle [BMNS13, BCB13, EKM^{+}18]. minds [BBS18]. Mining [LPGC20, Vau08]. mirrored [KK17a]. mixed [ZF12]. mobile [Ano22x, BJ15]. mode [Ano22p]. model [Ano22n, BVNS21]. modeling [MRSE21]. Modifications [AA17]. Modularity [WX14]. most [BTML16, BBS18]. Motivating [AOR14]. motivation [CZ20]. Moving [BS12, Tij08]. Mul’chenko [Rou21]. Multi [Mus14, SJM16]. multi-author [SJM16]. Multi-Authored [Mus14]. multidimensional [YBL17]. multidisciplinary [Ano22p]. Multiple [RS08, ABM19, Ano22k, LPGC20]. multivariate [YBL17]. Mutual [PNA15]. Mysore [Asu11, Asu11]. Myths [Glä08].

Nalimov [Rou21]. Nancy [KLJ17, RY18]. nanobiotechnology [dJSPR21].

Nanomaterials [RM09]. Nanosafety [LGJ14]. nanotechnology [Dar12, MES15, BS12, KGB11, GW21]. National [MLMJ13, MK12, SA07, WKK10, DSM20, MS13, VB17]. Nature [Ano22r]. Naukometriya [Rou21]. necessity [ABM19]. Network [GRAY15, KK07b, iKhJ14, LC15, Neg21, OKRS14, PPB09, dSVS12, WX14, Ano22h, Dar12, MES15, BM21, OB10, OCK17]. networking [RN17]. Networks [CF13, HLK^{+}09, HWL^{+}13, EBS16, KK17a, SJM16, Sun07]. Neural [PPB09]. Neuroscience [SS14]. News [AO07]. Nigeria [UE19]. NIRF [KTCA19]. NITs [VB17]. non [Kok19]. non-funded [Kok19]. normalization [Zit13]. normalized [MSH08]. North [NL16]. northeast [VB17]. Novel [Ano22n]. November [KJ16]. Number [Glä11, Sci14, SNK11, Ano22r, Rao12].

Observations [Hér08, VR15]. obtain [Glä11]. Obtained [KKS14]. Occurrence [TKSB18, RSM^{+}19]. Occurrences [MT08]. October [Asu11]. Odisha [GK16]. old [MS07b]. one [ZF12, Ahm18, SR17]. online [EE20]. Open [RC13, Ano22b, GR21]. Opening [dB08]. Operation [GRAY15]. Operations [AA09]. Organic [ES20, Ano22v, MFH18]. Organization [CKM16, HNN09]. Osareh [Osa08]. OSDI [ES16a]. other [KGBN12]. outcome [VT19]. Outcomes [RR14]. Outgrow [RR14]. Output [Bas14, Gup10, GBK11, GB11, GBG13, GG14, HS15, LMvI08, MS14, dSVS12, Ban19, DGG18, AE21, GK19, MG20, GD19b, KMG11, LLZ16, MLA12, Shu19]. outputs [BMNS13, OKRS14]. Ovary [Ram14]. Overall [PG09]. Overview [Cho12, Iva08, SCdM^{+}12]. Özel [Öze11].

Pair [GKL08, KK07a, KK08, PKL09, WKK10, KdBKK08]. Pairs [KK08, KdBKK08]. Pakistani [WWT21]. Paper [GP15, KZZV20, ZKJK19]. Papers [Ano18, Ano19a, Ano19b, Ano22a, KN11, Mus14, SGP16, Ano22k, HH20, Ho18, Naz18, YS19, YS20, ZKZ19]. Parallel [GD19b]. Part [KK08, KdBKK08]. past [ACO21, ES20]. Patent [CF13, KhJ13, YLX14, LS19]. Patenting [BP10, LS19]. path [LPGC20]. Pattern [ACO21, SM09, ZS12, GK19, SP20, SMM08]. Patterns [ES16b, KKS14, OB10]. peer [BK17]. Pen [Agu08b, Cha07b, Mey07b]. Percentile [PG09, BH17]. percentiles [OAdS21]. Performance [AAFH14, EGJ13, LC15, MLJM14, Muk11, PG09, Ano22e, Ano22v, MLMJ13, OP10, SM18, WH19]. Period [Hér08, OCK17]. peripheral [FPK21]. Personal [GRAY15]. Personality [VT19]. Personalization [Har14]. Perspective [Ram14, Ano22t, Ano22v, BJ15, LFH18, MES15, Saf14]. perspectives [Tij08]. petrochemical [MH17]. Ph.D. [SK11, SB14]. Pharmaceutics [KG09]. Pharmacology [KG09]. Philosophy [Hér08]. Phosphoinositide [HH20]. Physical [WKK10]. Physics [SK11, Sud10]. PI3K [HH20]. Planning [Ali10]. plant [Dis15]. Platform [JKK13, MS12a]. PLOS [Ahm18, SR17]. plus [TKSB18]. PNAS [KKS14]. Poisson [PB18]. Poland [MS13, OP10]. policy [HYWH20]. Polish [MS12b]. political [ÖP12]. Polycystic [Ram14]. Polymer [PG11]. Popular [YLZS15]. Portrait [SS10]. Positions [LMvI08]. posted [ÖP12]. Posthumous [Glä11]. postponement [Mar20]. potential [RVF^{+}13]. Power [iKhJ14]. Predicted [Ano22o]. Predicting [Ano22p, Ano22n]. Predictors [Pea20]. Preliminary [Hér12, TKSB18]. Presence [AMN17, JBM10]. present [ACO21]. presented [MS12b]. Presidents [GDY14]. Prevalence [KK17b]. principle [Ano22k]. problem [Zit13]. Process [Bur09]. Processes [AA09, Egg13]. processing [Ano22b]. procurement [RP21]. Production [AT15, Egg13, SCdM^{+}12]. productions [EKM^{+}18, RVF^{+}13]. Productivity [EGK08, EJHG13, GDY14, cMhLyW14, HB21, DKK18, LM21, PG11, WZ17]. Prof. [Hus09, Osa08, Zey09]. Professionals [MS12a]. Professor [KK12, San10, Sha10]. Profile [DKG10, Ram14, GK16, SMA15]. Program [Ron11, RR14]. Programme [Mus07]. projects [Mus17, RN17]. Promoting [MG12]. proposed [Ano22e]. provincial [LS19]. provincial-level [LS19]. Psychiatry [EGJ13]. Psychology [BCB13, EGJ13, VT19]. Public [MG12, Saf14, RP21]. Publication [BTML16, Bur09, DKK18, Gup10, KGB11, KMA20, Ano22u, Ban19, MLMJ13, NKK15, OB10, Shu19, ES20, YBL17]. Publication/Citation [Bur09]. Publications [DPR^{+}18, GBK11, GB11, GG14, MMM09, Ano22m, ACO21, DGG18, GD19b, HYWH20, JWH20, KTCA19, LM21, MIS^{+}20, SP20, SWH19, Ver09, HW20]. Published [Sin15, Ho18]. PubMed [Ano22q, MIS^{+}20, SMA15]. pure [kWhHR07]. pylori [SWH19].

Qualitative [Hér12, ZF12]. Quality [DSM20, MMM09, ASW^{+}19, MRAA20]. Quantitative [GBG13, cMhLyW14, RRB17, SG10, Glä08, SJMN16, ZF12]. Quantity [MMM09]. Quantum [Ano22n]. Quartile [Adi19]. Queries [MG12].

Race [BS12]. Random [AA09]. Ranganathan [SS07]. Rank [Ela20, PG09, AAMH17, SSKM21]. Ranked [ME12, KTCA19]. Ranking [BWC08, BSdMAM15, HM10, MT08, VB17, ASW^{+}19, Ano22l, FPK21, MW12, MSH08, SJM16, RS18]. Rankings [AM10, ABM19, Ano22l, Ela20]. rate [BK17]. rational [HAN14]. Re [cSL14, YLX14]. Re-applicability [YLX14]. Re-examination [cSL14]. received [Bar15]. recognized [RY18]. Reference [BTML16, KMA20]. References [Rao12, Sci14, Sud10]. reflected [Gup10, JMS19, Ver09]. Reflections [dB13, dCPFA10, RL08]. region [VB17]. Regional [SM18, OP10]. relation [Kok19, NKK15]. Relations [Rao12, Sci14]. Relationship [AT15, Ano22r, HA15]. Relative [LMvI08, Ano22k, Ela20]. release [BSdMAM15]. Relevance [iKhJ14]. Relying [RVB^{+}14]. remarkable [KZZV20]. Remarks [dB08]. Report [Asu11, KJ16, KLJ17, MK15, MKP19, YLM20]. repository [EGK08, GR21]. representativeness [AG17]. Represented [HM10]. Requirements [Ali11]. Research [Ano22s, dB13, BJ15, Cho12, DKG10, AE21, ES16b, GRAY15, GBK11, GB11, GBG13, GG14, Har14, HS15, KGBN12, KMG11, LGJ14, LLZ16, cSL14, cMhLyW14, MLMJ13, MLJM14, MJC19, LM21, MS14, Muk11, Mus11, RM09, RR15a, Ram14, RS11, Ron11, SS14, SK17, SM09, SA07, SM18, SG10, SKNN08, SNK11, Tij08, Vau08, VR15, WDG07, GW21, Ano22m, Ano22n, Ano22q, Ano22v, Ano22w, HB21, Bha17, BSdMAM15, BTML16, BK17, DKK18, Dis15, DK17, EBS16, EE20, FPK21, MG20, GD19a, GD19b, HH20, JWH20, JLH21, JSM20, JMS19, Kok19, KZZV20, KTCA19, LFH18, MG18, MLA12, MFH18, MRSE21, MP17, OM08, SP20, Pea20, RN17, Saf14, Shu19, Sun07, YS20, TKSB18, VT19, VHA21, WWH20, WZ17, CZ20, HAM^{+}21]. research-based [FPK21]. research-focused [BSdMAM15]. Researcher [DM14, OKRS14, OCK17, Rah20, RVF^{+}13]. Researchers [AAFH14, BBS18, EGJ13, Ano22e, SB14]. ResearchRabbit [Ano22s]. Resources [GD19a]. Response [MG12]. Result [MT08]. results [Ano22l]. Retrieval [LR14]. retrospect [LOA16]. Review [PNA15, ASW^{+}19, Ano22c, Ano22e, Ano22w, Ano22x, HAM^{+}21, HNN09, MHS16, SWH19, WZ17, ZKJK19, SDK14]. reviews [WH19]. Revisited [Rao12, Sud10]. rise [SSKM21]. Role [GDY14, PNA15, RVB^{+}14]. RPYS [BTML16, KMA20]. RRI [Ela20]. rules [DSM20]. Russia [MLA12, MS13, ES20]. Russian [AG17, AM10, MLJM14].

S [San10, Sha10, LG09, Ver09]. SAARC [SM18]. Sangam [San10]. SARS [Ano22q]. Scattering [Sud10, BVNS21]. scenarios [dJSPR21]. Scholar [LOA16, ZS12]. Scholarly [RC13, CKJC13]. Scholars [AST^{+}12, AMN17, GK19, LOA16]. Scholarship [Hér12]. Science [CWLL11, CKJC13, Gup10, HS15, HNN09, HAN14, HA15, HM10, KK17b, NK20, LC15, MR16b, OM11, OKRS14, PKL09, SMM08, SM14, SM18, SCdM^{+}12, UE19, Ver09, WDG07, ZKZ19, ASW^{+}19, Ano22h, BVNS21, DK17, Glä08, Ho18, KK17a, MW12, LM21, MRAA20, Nam17, OP10, OCK17, RRB17, Saf14, SJMN16, MS21, YS19, ZF12, BMNS13, HB21, dCPFA10, HS15, HM10, JMS19, LLO09, MW12, MJC19, MS14, OCK17, SWH19, WH19, Kal16, KKS14]. ScienceNet [YLZS15]. Sciences [AOR14, CF13, Hér08, KA08, Muk11, SM09, AG17, Ano22d, Ano22p, MS07b, RN17, Rou12, GR21, TKSB18, MLJM14, UE19]. Scientific [AAFH14, BMNS13, dB13, BCB13, ES15, EJHG13, HWL^{+}13, LMvI08, MMM09, PG11, RRA14, dJSdJP21, dSVS12, AWAMA18, AOR14, BBS18, BH12, EKM^{+}18, JAAN14, MS07b, MSH08, MS12b, MPPH17, MIS^{+}20, Neg21, NK14, NKK15, OAR14, OKRS14, Rah20, RRB17, RVF^{+}13, SA16, Sha19, WZ17, Zit13]. scientist [BH12, KKYR17]. Scientists [Bas14, Ahm18, Ano22g, DKK18, Ho18, Naz18, PG11, SK17]. Scientometric [Ano22t, Ano22u, Ano22v, Ano22w, HB21, BS12, DKG10, EGJ13, GK16, GDY14, GB11, GG14, HM10, KGBN12, KG09, KMG11, KA08, MG18, MP17, SS14, SMM08, SS10, VR15, WDG07, Ahm18, DGG18, Dis15, GD19a, GD19b, LLZ16, MJC19, Pal16, SP20, SK17, SS19, SM18, Shu19, WWT21, WZ17, CZ20]. Scientometrical [dCPFA10]. Scientometrics [Ano18, Ano19a, Ano19b, Ano22a, Asu11, GB21, DHZK09, GBK11, HLK^{+}09, HWL^{+}13, Iva08, KN11, KJ16, KLJ17, MK15, MKP19, Mar20, PNA15, RR15b, SAPM13, SMA15, YLM20, ZF12, GK19, RSM^{+}19, SA16, EGK08, RS08, MR19]. SCImago [RS18, AAMH17]. SciSearch [OM08]. SCOPUS [MJC19, DPR^{+}18, Dis15, EJHG13, MP17, MR16b, OAdS21, RSZ15, Shu19, VT19, ZKZ19]. score [AAMH17]. scores [YBL17]. Search [MT08, MG12, EE20, ME12]. Secondary [hJjJiK13]. Sector [BP10, MLJM14, PRP21]. seen [SMA15]. Selected [JBM10, ME12]. self [ÖP12]. self-images [ÖP12]. Semi [MT08]. Semi-Automated [MT08]. Semiconductor [WX14]. seminal [BTML16, KMA20]. service [OCK17]. services [HYWH20]. Sets [MT08]. Seven [Glä08]. Shafi [Sha10]. Shahid [AOR14]. Shapes [KK08]. Share [GRAY15, BH12]. sharing [ZKJK19]. SIC [Hér08]. side [Zit13]. Signatures [LR14]. Similar [RR14]. simple [RR14]. Sites [Vau12]. Size [CWLL11, NKK15]. Sketch [Agu08a, Cha07a, Hus09, Mey07a, Osa08, Öze11, San10, Sha10, Zey09]. skyline [SSKM21]. sleeping [KZZV20, ZKZ19]. Small [Mus14]. Social [KK07a, KK07b, LC15, SM14, AG17, Ano22d, Dar12, MS07b, TKSB18, OKRS14]. Societies [LLO09]. Software [BTML16, RR14, ZKZ19, Kok19, MS21]. Solutions [OAR14]. some [BBS18, Glä11, Tij08]. SOSP [ES16a]. Source [Vau12]. sources [Ano22d, UY17]. South [Gup10, SM18, SCdM^{+}12]. Space [MT09, Hol11, SK17]. special [Rou12]. Specialization [Hér12]. Spectroscopy [BTML16, KMA20]. spectrum [KMA20]. speech [RN17]. Spices [SA07]. Springer [Ano22r]. squaraine [Ano22w]. Sri [Dis15]. SSCI [OKRS14, MS07a]. state [GK16]. States [LM21, MES15]. Statistical [VT19, RS18]. Statistics [JKK13]. Status [Gup10]. Stem [KGBN12, dJSdJP21]. STI [Nam17]. STM [JKK13]. storm [SP20]. Strategic [Ali11]. strategies [dJSPR21]. Strongly [GP15]. Structural [RR14]. Structure [KKS14, OKRS14, OM08]. students [UY17]. Studies [KK08, KdBKK08, Ano22h, Ano22j, Glä08, GD19a, MS21, NK20]. Study [BP10, BS12, BCB13, EBS16, EGJ13, HM10, JBM10, KGB11, KS10, KA08, LG09, MS14, RT15, RS08, RVB^{+}14, SGP16, Sin15, WDG07, YLC13, ZKJK19, Ano22q, AO07, GB21, Bur09, DKK18, Dis15, DSM20, AE21, GK16, HS15, Ho18, JAAN14, KGBN12, KMG11, KTCA19, NK20, LLZ16, LBN21, MS13, MJC19, MH17, MP17, Mus17, MR16b, NL16, NK14, SP20, PG11, RS11, Saf14, SK17, STP18, SS19, Shu19, Sud10, SR17, TKSB18, UE19, Ver09]. Subject [Hér12, Ano22r, YS19]. Submitted [SK11]. Success [BK17]. Summary [DHZK09]. Sums [AA09]. Supercomputing [DGG18]. Superiority [PG09]. surges [SP20]. sustainability [Mus17]. sustainable [YS19]. Syndrome [Ram14]. System [Ali10, Ali11, ES16a, KMA20, RSM^{+}19]. Systematic [Ano22x]. Systems [GRAY15].

T [LG09, Ver09]. table [RY18]. Taiwanese [Ho18]. taking [kWhHR07]. tamer [RY18]. Teaching [KK12]. Teamwork [dB13]. Technical [hJjJiK13, KhJ13, MS21]. technique [OM08]. Technological [dJSPR21]. Technologies [cSL14, Sin15]. Technology [BP10, DV18, Gup10, Har14, HM10, iKhJ14, PKL09, VB17, Ano22x, Nam17, PRP21, SJMN16, SK17, YS19, VR15, GKL08, UE19]. Technovation [LLZ16]. Tehran [Saf14]. Ten [ES15, HAM^{+}21]. Their [ES16b, AM10, MLMJ13]. theory [Ano22x, MG18, MJC19]. Theses [SK11, SB14, Sud10, UY17]. third [BSdMAM15]. Thirteenth [KN11]. Thirty [MHS16]. thoughts [Glä11]. three [Ano22r, JSM20]. time [OCK17]. Today [Iva08]. tools [MS21]. Top [ES15, EH18, EJHG13, ME12, SGP16, AE21, KTCA19, OAdS21, YS19, YS20]. Top-cited [EH18]. Topic [MT08, RM09, MRSE21]. Toxicology [KG09]. traits [UY17]. trajectories [LPGC20]. Transfer [RVB^{+}14]. Transition [Hér08]. travel [CZ20]. Trend [EKM^{+}18, KGB11, MJC19]. Trends [PK15, VR15, Ano22u, BJ15, DV18, DK17, JWH20, Lev07, SP20, Sun07]. tribology [EBS16, EH18]. trust [ZKJK19]. Tuberculosis [GB11]. Turkey [Dar12]. Turkish [AST^{+}12]. twenty [HB21]. Two [MLJM14, RS08, MS21]. type [Ano22k, RL08, YBL17]. types [Ano22c].

‘UHasselt’ [EGK08]. Ukrainian [Naz18]. ulcer [Ban19]. under-recognized [RY18]. Unified [Ano22x]. United [LM21, MES15]. Universities [BWC08, ES15, GDY14, HM10, KTCA19, MLMJ13, Saf14, SGP16, BSdMAM15, SSKM21, Sun07]. University [Asu11, DPR^{+}18, KJ16, MKP19, OB10, SNK11, ABM19, FPK21, LS19, AOR14, AM10, DSM20, MS14, NL16, SK11]. Updated [dB13]. URLs [KK17b]. USA [LMvI08]. Usability [MK12, UY17]. Usage [BWC08, JKK13, Vau08]. Use [JKK13, Ano22s, Ano22x, dJSPR21]. used [SB14]. Using [PPB09, ZS12, Ano22p, Ano22s, AO07, DSM20, iKhJ14, LR14, LPGC20, LFH18, LLO09, MRAA20, MRSE21, BM21, OB10, OM08, OM11, ÖP12, RSM^{+}19, Shu19, SG10, VT19]. UTAUT [Ano22x]. Utilising [AA09]. Utility [WX14]. Utilizing [JKK13].

vaccine [Ano22q]. Valid [SS07]. validity [BK17]. values [BH17]. Variability [Zit13]. Varying [KK08]. Vector [RVB^{+}14]. vectors [YBL17]. versus [SA16]. via [MG12, SJMN16]. Virtual [SS07]. virus [dJSPR21, Ano22q]. Visibility [LOA16, NK14]. visualisation [Ano22t]. visualization [Ano22s, MS07b, STP18, GW21, DHZK09]. Visualizing [OM08, RR14]. volatile [MFH18]. Volume [Ano11, Ano12, Ano22n].

warming [SS19]. Weak [Sci14]. Web [dCPFA10, Rou12, NK14, VB17, BMNS13, HB21, BWC08, DV18, GRAY15, HS15, Hol11, JBM10, JMS19, NK20, MW12, MJC19, MG12, MK12, OM11, OCK17, PK15, RC13, Sin15, Vau08, Vau12, WH19, ZKZ19]. webometric [LLO09, JAAN14, JBM10, KK07b, ME12, GR21, Vau08]. Webometrics [Ano18, Ano19a, Ano19b, Ano22a, Asu11, DHZK09, GRAY15, KN11, KJ16, KLJ17, MK15, MKP19, Mar20, YLM20, AO07]. website [DV18, MK12]. Websites [LR14, AO07, JAAN14, LLO09]. Weighted [AAFH14, Adi19]. West [RS11]. Wikipedia [NKK15]. Wind [iKhJ14]. WIS [Ano18, Ano19a, Ano19b, Ano22a, KN11, KJ16, KLJ17, MK15, MKP19, Mar20, YLM20]. within [ME12]. women [GK19]. word [ÖP12]. work [OAR14]. works [BTML16, KMA20]. World [GBG13, MS13, BBS18, MSH08, MS12b]. worldwide [BSdMAM15]. WOS [BMNS13, OM11]. writing [Pea20]. written [Naz18, Pea20]. www.slp.org.in [KJ16].

Year [BTML16, KMA20, MR16a]. Year-Based [MR16a]. Years [KK12, HB21, HAM^{+}21, MHS16, ES20]. yoga [MIS^{+}20].

Zeyuan [CWS20, KK12, Zey09]. Zika [dJSPR21].

References

Artikis:2009:LPI

[AA09] Constantinos T. Artikis and Panagiotis T. Artikis. Learning processes and information management operations utilising a class of random sums. Collnet Journal of Scientometrics and Information Management, 3(1):31–38, 2009. CODEN ???? ISSN 0973-7766 (print), 2168-930X (electronic).

Alguliyev:2017:MJI

[AA17] Rasim M. Alguliyev and Ramiz M. Aliguliyev. Modifications to the journal impact factor. Collnet Journal of Scientometrics and Information Management, 11(1):25–43, 2017. CODEN ???? ISSN 0973-7766 (print), 2168-930X (elec-

            tronic). URL http://www.tandfonline.com//doi/abs/10.
            1080/09737766.2016.1235251.

Alguliev:2014:WCI

[AAFH14] Rasim M. Alguliev, Ramiz M. Aliguliyev, Tahmasib Kh. Fataliyev, and Rahila Sh. Hasanova. Weighted consensus index for assessment of the scientific performance of researchers. Collnet Journal of Scientometrics and Information Management, 8 (2):371–400, 2014. CODEN ???? ISSN 0973-7766 (print), 2168-930X (electronic).

Ahmad:2017:CAJ

[AAMH17] Shakil Ahmad Jamil Ahmad, Isam Mohammed Abdel-Magid, and Akhtar Hussain. Comparison among journal impact factor, SCimago journal rank indicator, eigenfactor score and h5index of environmental engineering journals. Collnet Journal of Scientometrics and Information Management, 11(1):133–151, 2017. CODEN ???? ISSN 0973-7766 (print), 2168-930X

(electronic). URL http://www.tandfonline.com//doi/abs/
10.1080/09737766.2016.1266807.

Angelis:2019:NMU

[ABM19] Lefteris Angelis, Nick Bassiliades, and Yannis Manolopoulos. On the necessity of multiple university rankings. Collnet Journal of Scientometrics and Information Management, 13(1):11– 36, 2019. CODEN ???? ISSN 0973-7766 (print), 2168-930X

(electronic). URL http://www.tandfonline.com/doi/abs/10.
1080/09737766.2018.1550043.

Ayaz:2021:PDC

[ACO21] Ahmet Ayaz, Kamil Celik, and Ozcan Ozyurt. Pattern detection in cloud computing: Bibliometric mapping of publications in the field from past to present. Collnet Journal of Scientometrics and Information Management, 15(2):469–494, 2021. CODEN ???? ISSN 0973-7766 (print), 2168-930X

(electronic). URL http://www.tandfonline.com/doi/abs/10.
1080/09737766.2021.2007038.

Adigozalova:2019:QWI

[Adi19] Narmin A. Adigozalova. Quartile weighted impact factor. Collnet Journal of Scientometrics and Information Management, 13(2): 365–386, 2019. CODEN ???? ISSN 0973-7766 (print), 2168-930X

(electronic). URL http://www.tandfonline.com/doi/abs/10.
1080/09737766.2020.1716646.

Elbadawy:2021:ROT

[AE21] Waleed Ali and Ahmed Elbadawy. Research output of the top 10 African countries: an analytical study. Collnet Journal of Scientometrics and Information Management, 15(1):9– 25, 2021. CODEN ???? ISSN 0973-7766 (print), 2168-930X

(electronic). URL http://www.tandfonline.com/doi/abs/10.
1080/09737766.2021.1934181.

Anisimova:2017:RRS

[AG17] A. E. Anisimova and E. G. Grebenshchikova. The representativeness of the Russian social sciences and the humanities in the global information flows. Collnet Journal of Scientometrics and Information Management, 11(2):299–307, 2017. CODEN ???? ISSN 0973-7766 (print), 2168-930X (elec-

tronic). URL http://www.tandfonline.com//doi/abs/10.
1080/09737766.2017.1306180.

Aguillo:2008:BSD

[Agu08a] Isidro F. Aguillo. Biographical sketch of Dr. Isidro F. Aguillo. Collnet Journal of Scientometrics and Information Management, 2(1):v, 2008. CODEN ???? ISSN 0973-7766 (print), 2168-930X (electronic).

Aguillo:2008:DAP

[Agu08b] Isidro F. Aguillo. From Dr. Aguillo’s pen. Collnet Journal of Scientometrics and Information Management, 2(1):vii, 2008. CO-DEN ???? ISSN 0973-7766 (print), 2168-930X (electronic).

Ahmadi:2018:CIS

[Ahm18] Ashiya Ahmadi. Contribution of Indian scientists in PLOS ONE: a scientometric analysis. Collnet Journal of Scientometrics and Information Management, 12(2):183–196, 2018. CODEN ???? ISSN 0973-7766 (print), 2168-930X (elec-

tronic). URL http://www.tandfonline.com/doi/abs/10.
1080/09737766.2018.1433101.

Ali:2010:CPH

[Ali10] S. M. Wazed Ali. Concept of planning: Higher education management information system. Collnet Journal of Scientometrics and Information Management, 4(2):13–19, 2010. CODEN ???? ISSN 0973-7766 (print), 2168-930X (electronic).

Ali:2011:SRH

[Ali11] S. M. Wazed Ali. Strategic requirements of higher education management information system. Collnet Journal of Scientometrics and Information Management, 5(2):145–155, 2011. CODEN ???? ISSN 0973-7766 (print), 2168-930X (electronic).

Aversa:2010:BIT

[AM10] E. Aversa and V. Markusova. Bibliometric indicators and their impact on Russian University rankings. Collnet Journal of Scientometrics and Information Management, 4(1):1–8, 2010. CODEN ???? ISSN 0973-7766 (print), 2168-930X (electronic).

Asnafi:2017:PIS

[AMN17] Amir Reza Asnafi, Shima Moradi, and Maryam Pakdaman Naeini. Presence of Iranian scholars in COLLNET conferences (2003–2015). Collnet Journal of Scientometrics and Information Management, 11(2):193–200, 2017. CODEN ???? ISSN 0973-7766 (print), 2168-930X (electronic). URL http:

//www.tandfonline.com//doi/abs/10.1080/09737766.2017.
1284728.

Anonymous:2007:FMa

[Ano07a] Anonymous. Front matter. Collnet Journal of Scientometrics and Information Management, 1(1):i, 2007. CODEN ???? ISSN 0973-7766 (print), 2168-930X (electronic).

Anonymous:2007:FMb

[Ano07b] Anonymous. Front matter. Collnet Journal of Scientometrics and Information Management, 1(2):i, 2007. CODEN ???? ISSN 0973-7766 (print), 2168-930X (electronic).

Anonymous:2011:IV

[Ano11] Anonymous. Index of volume 5 (2011). Collnet Journal of Scientometrics and Information Management, 5(2):243–244, 2011. CODEN ???? ISSN 0973-7766 (print), 2168-930X (electronic).

Anonymous:2012:IV

[Ano12] Anonymous. Index of volume 6 (2012). Collnet Journal of Scientometrics and Information Management, 6(2):335–337, 2012. CODEN ???? ISSN 0973-7766 (print), 2168-930X (electronic).

Anonymous:2018:ACP

[Ano18] Anonymous. Announcement and call for papers 14^{th} International Conference on Webometrics, Informetrics and Scientometrics (WIS) and 19^{th} COLLNET Meeting. Collnet Journal of Scientometrics and Information Management, 12(1):161– 164, 2018. CODEN ???? ISSN 0973-7766 (print), 2168-930X

(electronic). URL http://www.tandfonline.com//doi/abs/
10.1080/09737766.2018.1488657.

Anonymous:2019:ACPa

[Ano19a] Anonymous. Announcement & call for papers 15^{th} International Conference on Webometrics, Informetrics and Scientometrics (WIS) & 20^{th} COLLNET Meeting. Collnet Journal of Scientometrics and Information Management, 13(1):7– 10, 2019. CODEN ???? ISSN 0973-7766 (print), 2168-930X

(electronic). URL http://www.tandfonline.com/doi/abs/10.
1080/09737766.2019.1617259.

Anonymous:2019:ACPb

[Ano19b] Anonymous. Announcement and call for papers: 16^{th} International Conference on Webometrics, Informetrics and Scientometrics (WIS) and 20^{th} COLLNET Meeting. Collnet Journal of Scientometrics and Information Management, 13(2):213– 217, 2019. CODEN ???? ISSN 0973-7766 (print), 2168-930X

(electronic). URL http://www.tandfonline.com/doi/abs/10.
1080/09737766.2019.1746513.

Anonymous:2022:ACP

[Ano22a] Anonymous. Announcement and call for papers 16^{th} International Conference on Webometrics, Informetrics and Scientometrics (WIS) and 21^{st} COLLNET Meeting. Collnet Journal of Scientometrics and Information Management, 16(1):3–6, 2022. CODEN ???? ISSN 0973-7766 (print), 2168-930X (electronic).

Anonymous:2022:IMA

[Ano22b] Anonymous. Are indexing, metrics and the article processing charges (APC) of Hindawi open access journals linked? Collnet Journal of Scientometrics and Information Management, 16(2): 239–251, 2022. CODEN ???? ISSN 0973-7766 (print), 2168-930X (electronic).

Anonymous:2022:BCT

[Ano22c] Anonymous. Bibliographic coupling and types of centralities: a review of ASME journals — 2000 to 2020. Collnet Journal of Scientometrics and Information Management, 16(1):161–186, 2022. CODEN ???? ISSN 0973-7766 (print), 2168-930X (electronic).

Anonymous:2022:BMS

[Ano22d] Anonymous. Bibliometric measures in social sciences and humanities from different sources based on Brazilian data. Collnet Journal of Scientometrics and Information Management, 16(2):

279–304, 2022. CODEN ???? ISSN 0973-7766 (print), 2168-930X (electronic).

Anonymous:2022:CRP

[Ano22e] Anonymous. A critical review of the proposed academic performance indicators for the assessment of individual researchers in Hungary. Collnet Journal of Scientometrics and Information Management, 16(2):331–352, 2022. CODEN ???? ISSN 0973-7766 (print), 2168-930X (electronic).

Anonymous:2022:DIG

[Ano22f] Anonymous. The development of Indonesian e-government: a bibliometric analysis. Collnet Journal of Scientometrics and Information Management, 16(1):49–74, 2022. CODEN ???? ISSN 0973-7766 (print), 2168-930X (electronic).

Anonymous:2022:GIN

[Ano22g] Anonymous. GH -index: a new index for the assessment of scientists. Collnet Journal of Scientometrics and Information Management, 16(2):407–464, 2022. CODEN ???? ISSN 0973-7766 (print), 2168-930X (electronic).

Anonymous:2022:ECS

[Ano22h] Anonymous. Examination of e-commerce studies between 2016 and 2021 with science mapping and network analysis. Collnet Journal of Scientometrics and Information Management, 16(2): 477–507, 2022. CODEN ???? ISSN 0973-7766 (print), 2168-930X (electronic).

Anonymous:2022:ED

[Ano22i] Anonymous. From the Editor’s desk. Collnet Journal of Scientometrics and Information Management, 16(1):1–2, 2022. CODEN ???? ISSN 0973-7766 (print), 2168-930X (electronic).

Anonymous:2022:GDD

[Ano22j] Anonymous. Gender differences, data carpentry and bibliometric studies in mathematics. Collnet Journal of Scientometrics and Information Management, 16(2):465–476, 2022. CODEN ???? ISSN 0973-7766 (print), 2168-930X (electronic).

Anonymous:2022:TIM

[Ano22k] Anonymous. h-type indices for multiple authorship papers based on “relative first author” principle. Collnet Journal of Scientometrics and Information Management, 16(2):305–330, 2022. CO-DEN ???? ISSN 0973-7766 (print), 2168-930X (electronic).

Anonymous:2022:HGR

[Ano22l] Anonymous. Hegemony in global rankings: a Gramscian analysis of bibliometric indices and ranking results. Collnet Journal of Scientometrics and Information Management, 16(2):253–277, 2022. CODEN ???? ISSN 0973-7766 (print), 2168-930X (electronic).

Anonymous:2022:MIC

[Ano22m] Anonymous. Measuring the impact of co-author count on citation count of research publications. Collnet Journal of Scientometrics and Information Management, 16(1):35–48, 2022. CODEN ???? ISSN 0973-7766 (print), 2168-930X (electronic).

Anonymous:2022:NMP

[Ano22n] Anonymous. Novel model for predicting the future volume of research articles on applications of quantum dots. Collnet Journal of Scientometrics and Information Management, 16(1):187–213, 2022. CODEN ???? ISSN 0973-7766 (print), 2168-930X (electronic).

Anonymous:2022:PC

[Ano22o] Anonymous. Predicted correlation. Collnet Journal of Scientometrics and Information Management, 16(1):7–18, 2022. CO-DEN ???? ISSN 0973-7766 (print), 2168-930X (electronic).

Anonymous:2022:PAM

[Ano22p] Anonymous. Predicting access mode of multidisciplinary and library and information sciences journals using machine learning. Collnet Journal of Scientometrics and Information Management, 16(1):117–124, 2022. CODEN ???? ISSN 0973-7766 (print), 2168-930X (electronic).

Anonymous:2022:PBB

[Ano22q] Anonymous. A PubMed based bibliometric study on Covid-19 and SARS Corona Virus vaccine and alternative medicine-based research and development. Collnet Journal of Scientometrics and Information Management, 16(2):389–405, 2022. CODEN ???? ISSN 0973-7766 (print), 2168-930X (electronic).

Anonymous:2022:RBN

[Ano22r] Anonymous. Relationship between number of downloads and three journal-based metrics of 11 subject categories among 1575 Springer Nature journals. Collnet Journal of Scientometrics and Information Management, 16(2):371–388, 2022. CODEN ???? ISSN 0973-7766 (print), 2168-930X (electronic).

Anonymous:2022:RDV

[Ano22s] Anonymous. Research discovery and visualization using ResearchRabbit: a use case of AI in libraries. Collnet Journal of Scientometrics and Information Management, 16(2):215–237, 2022. CODEN ???? ISSN 0973-7766 (print), 2168-930X (electronic).

Anonymous:2022:SAV

[Ano22t] Anonymous. Scientometric analysis and visualisation of global information literacy from higher education perspective. Collnet Journal of Scientometrics and Information Management, 16(1): 125–143, 2022. CODEN ???? ISSN 0973-7766 (print), 2168-930X (electronic).

Anonymous:2022:SMG

[Ano22u] Anonymous. Scientometric mapping of global publication trends in health informatics domain. Collnet Journal of Scientometrics and Information Management, 16(1):145–159, 2022. CO-DEN ???? ISSN 0973-7766 (print), 2168-930X (electronic).

Anonymous:2022:SMO

[Ano22v] Anonymous. Scientometric mapping of organic farming research performance in the global perspective during 2010–2019. Collnet Journal of Scientometrics and Information Management, 16(1): 19–33, 2022. CODEN ???? ISSN 0973-7766 (print), 2168-930X (electronic).

Anonymous:2022:SRG

[Ano22w] Anonymous. Scientometric review of global squaraine dye research. Collnet Journal of Scientometrics and Information Management, 16(2):353–370, 2022. CODEN ???? ISSN 0973-7766 (print), 2168-930X (electronic).

Anonymous:2022:UTA

[Ano22x] Anonymous. Unified theory of acceptance and use of technology (UTAUT) in mobile learning adoption : Systematic literature review and bibliometric analysis. Collnet Journal of Scientometrics and Information Management, 16(1):75–116, 2022. CODEN ???? ISSN 0973-7766 (print), 2168-930X (electronic).

Asnafi:2007:SCA

[AO07] Amir Reza Asnafi and Farideh Osareh. A study of collaboration among Iranian news agencies websites using Webometrics methods. Collnet Journal of Scientometrics and Information Management, 1(1):23–26, 2007. CODEN ???? ISSN 0973-7766 (print), 2168-930X (electronic).

Ahmadi:2014:MIF

[AOR14] Elham Ahmadi, Farideh Osareh, and Aref Riahi. Motivating and inhibiting factors of scientific collaboration of faculty members in Shahid Chamran University and Jundishapur Medical Sciences University in Ahvaz-Iran. Collnet Journal of Scientometrics and Information Management, 8(2):237–245, 2014. CO-DEN ???? ISSN 0973-7766 (print), 2168-930X (electronic).

Al:2012:CTS

[AST^{+}12] Umut Al, Irem Soydal, Zehra Taskin, Güleda Düzyol, and Umut Sezen. Collaboration of Turkish scholars: Local or global? Collnet Journal of Scientometrics and Information Management, 6 (1):145–159, 2012. CODEN ???? ISSN 0973-7766 (print), 2168-930X (electronic).

Asundi:2011:CRI

[Asu11] A. Y. Asundi. Conference report of the 6^{th} International Conference on Webometrics, Informetrics and Scientometrics and 11^{th} COLLNET Meeting, 19^{th}–22^{nd} October 2010. University of Mysore, Mysore. Collnet Journal of Scientometrics and Information Management, 5(1):129–143, 2011. CODEN ???? ISSN 0973-7766 (print), 2168-930X (electronic).

Ahmad:2019:EJQ

[ASW^{+}19] Shakil Ahmad, Md Sohail, Abu Waris, Isam Mohammed Abdel-Magid, Abdurahiman Pattukuthu, and Muhammed Shakir Azad. Evaluating journal quality: a review of journal citation indicators and ranking in library and information science core journals. Collnet Journal of Scientometrics and Information Management, 13(2):345–363, 2019. CODEN ???? ISSN 0973-7766 (print),

2168-930X (electronic). URL http://www.tandfonline.com/
doi/abs/10.1080/09737766.2020.1718030.

Al:2015:RBE

[AT15] Umut Al and Zehra Taskin. Relationship between economic development and intellectual production. Collnet Journal of Scientometrics and Information Management, 9(1):25–35, 2015. CO-DEN ???? ISSN 0973-7766 (print), 2168-930X (electronic).

Ahmad:2018:ASI

[AWAMA18] Shakil Ahmad, Abu Waris, Isam Mohammed Abdel-Magid, and Mohammad Aqil. An assessment of scientific impact of Indian journals. Collnet Journal of Scientometrics and Information Management, 12(1):133–147, 2018. CODEN ???? ISSN 0973-7766 (print), 2168-930X (electronic). URL http:

//www.tandfonline.com//doi/abs/10.1080/09737766.2018.
1433063.

Bansal:2019:BAG

[Ban19] Sonia Bansal. A bibliometric assessment of global publication output on Buruli ulcer during 2000–2017. Collnet Journal of Scientometrics and Information Management, 13(1):65– 77, 2019. CODEN ???? ISSN 0973-7766 (print), 2168-930X

(electronic). URL http://www.tandfonline.com/doi/abs/10.
1080/09737766.2018.1541043.

Barrantes:2015:ICR

[Bar15] Bárbara S. Lancho Barrantes. Increase in citations received by India from collaborating countries. Collnet Journal of Scientometrics and Information Management, 9(1):107–119, 2015. CO-DEN ???? ISSN 0973-7766 (print), 2168-930X (electronic).

Basu:2014:CIC

[Bas14] Aparna Basu. A comparison of international collaboration by Indian scientists and the diaspora output: Country and institutions. Collnet Journal of Scientometrics and Information Management, 8(1):169–182, 2014. CODEN ???? ISSN 0973-7766 (print), 2168-930X (electronic).

Bornmann:2018:HCR

[BBS18] Lutz Bornmann, Johann Bauer, and Elisabeth Maria Schlagberger. Highly cited researchers 2014 and 2015: An investigation of some of the world’s most influential scientific minds on the institutional and country level. Collnet Journal of Scientometrics and Information Management, 12(1):15– 33, 2018. CODEN ???? ISSN 0973-7766 (print), 2168-930X

(electronic). URL http://www.tandfonline.com//doi/abs/
10.1080/09737766.2017.1332610.

Biglu:2013:SSM

[BCB13] Mohammad-Hossein Biglu, Negin Chakhmachi, and Sahar Biglu. Scientific study of Middle East countries in psychology (1996– 2010). Collnet Journal of Scientometrics and Information Management, 7(2):293–296, 2013. CODEN ???? ISSN 0973-7766 (print), 2168-930X (electronic).

Bousari:2012:FAS

[BH12] Roghayeh Ghorbani Bousari and Mohammad Hassanzadeh. Factors that affect scientists’ behavior to share scientific knowledge. Collnet Journal of Scientometrics and Information Management, 6(2):215–227, 2012. CODEN ???? ISSN 0973-7766 (print), 2168-930X (electronic).

Bornmann:2017:EVP

[BH17] Lutz Bornmann and Robin Haunschild. Expected values in percentile indicators. Collnet Journal of Scientometrics and Information Management, 11(2):249–252, 2017. CODEN ???? ISSN 0973-7766 (print), 2168-930X (electronic). URL http:

//www.tandfonline.com//doi/abs/10.1080/09737766.2017.
1292668.

Bhardwaj:2017:MGR

[Bha17] Raj Kumar Bhardwaj. Mapping of global research literature in the field of cloud computing. Collnet Journal of Scientometrics and Information Management, 11(2):173–191, 2017. CODEN ???? ISSN 0973-7766 (print), 2168-930X (elec-

tronic). URL http://www.tandfonline.com//doi/abs/10.
1080/09737766.2017.1284727.

Bhardwaj:2015:RTM

[BJ15] Raj Kumar Bhardwaj and P. K. Jain. Research trends in mobile learning: A global perspective. Collnet Journal of Scientometrics and Information Management, 9(2):205–224, 2015. CODEN ???? ISSN 0973-7766 (print), 2168-930X

(electronic). URL http://www.tandfonline.com/doi/abs/10.
1080/09737766.2015.1069960.

Burhan:2017:SRV

[BK17] Ibrahim Burhan and Ho Sung Kim. Success rate validity of peer assessment for research grant applications. Collnet Journal of Scientometrics and Information Management, 11(2):341– 359, 2017. CODEN ???? ISSN 0973-7766 (print), 2168-930X

(electronic). URL http://www.tandfonline.com//doi/abs/
10.1080/09737766.2017.1321729.

Mota:2021:ECD

[BM21] Luiza Amara Maciel Braga and Fabio Batista Mota. Early cancer diagnosis using lab-on-a-chip devices: a bibliometric and network analysis. Collnet Journal of Scientometrics and Information Management, 15(1):163–196, 2021. CODEN ???? ISSN 0973-7766 (print), 2168-930X (electronic). URL http:

//www.tandfonline.com/doi/abs/10.1080/09737766.2021.
1949949.

Baji:2013:SOM

[BMNS13] Fatima Baji, Parastoo Parsaei Mohammadi, Zivar Sabaghi Nejad, and Somaye Sharify. Scientific outputs of Middle Eastern countries in Web of Science database (WOS) from 1990 to 2010. Collnet Journal of Scientometrics and Information Management, 7(1):111–120, 2013. CODEN ???? ISSN 0973-7766 (print), 2168-930X (electronic).

Bhattacharya:2010:ACD

[BP10] Sujit Bhattacharya and Swapan Kumar Patra. Assessing competency of a developing country in high technology?: A case study based on Indian patenting activity in the biotechnology sector. Collnet Journal of Scientometrics and Information Management, 4(2):21–34, 2010. CODEN ???? ISSN 0973-7766 (print), 2168-930X (electronic).

Bhattacharya:2012:CMA

[BS12] Sujit Bhattacharya and Shilpa. China moving ahead in the global Nanotechnology race: Evidences from scientometric study. Collnet Journal of Scientometrics and Information Management, 6 (1):97–117, 2012. CODEN ???? ISSN 0973-7766 (print), 2168-930X (electronic).

Bornmann:2015:RMU

[BSdMAM15] Lutz Bornmann, Moritz Stefaner, Felix de Moya Anegón, and Rüdiger Mutz. Ranking and mapping of universities and research-focused institutions worldwide: The third release of excellencemapping.net. Collnet Journal of Scientometrics and Information Management, 9(1):65–72, 2015. CODEN ???? ISSN 0973-7766 (print), 2168-930X (electronic).

Bornmann:2016:ISW

[BTML16] Lutz Bornmann, Andreas Thor, Werner Marx, and Loet Leydesdorff. Identifying seminal works most important for research fields: Software for the Reference Publication Year Spectroscopy (RPYS). Collnet Journal of Scientometrics and Information Management, 10(1):125–140, 2016. CODEN ???? ISSN 0973-7766 (print), 2168-930X (electronic).

Burrell:2009:PCP

[Bur09] Quentin L. Burrell. The publication/citation process at the micro level: A case study. Collnet Journal of Scientometrics and Information Management, 3(1):71–77, 2009. CODEN ???? ISSN 0973-7766 (print), 2168-930X (electronic).

Borgohain:2021:ABL

[BVNS21] Dhruba Jyoti Borgohain, Manoj Kumar Verma, Mohammad Nazim, and Mayukh Sarkar. Application of Bradford’s law of scattering and Leimkuhler model to information science literature. Collnet Journal of Scientometrics and Information Management, 15(1):197–212, 2021. CODEN ???? ISSN 0973-7766 (print), 2168-930X (electronic). URL http:

//www.tandfonline.com/doi/abs/10.1080/09737766.2021.
1943041.

Boell:2008:UDW

[BWC08] Sebastian K. Boell, Concepción S. Wilson, and Fletcher T. H. Cole. Usage of different Web impact factors for ranking Australian universities. Collnet Journal of Scientometrics and Information Management, 2(2):57–70, 2008. CODEN ???? ISSN 0973-7766 (print), 2168-930X (electronic).

Chen:2013:IEP

[CF13] Yunwei Chen and Shu Fang. Identifying the evolution of patent assignees collaboration networks: A case of Chinese Academy of Sciences (CAS). Collnet Journal of Scientometrics and Information Management, 7(1):69–80, 2013. CODEN ???? ISSN 0973-7766 (print), 2168-930X (electronic).

Chandra:2007:BSD

[Cha07a] Harish Chandra. Biographical sketch of Dr. Harish Chandra. Collnet Journal of Scientometrics and Information Management, 1(2):vii, 2007. CODEN ???? ISSN 0973-7766 (print), 2168-930X (electronic).

Chandra:2007:DCP

[Cha07b] Harish Chandra. From Dr. Chandra’s pen . . .. Collnet Journal of Scientometrics and Information Management, 1(2):ix, 2007. CODEN ???? ISSN 0973-7766 (print), 2168-930X (electronic).

Choi:2012:OIR

[Cho12] Heeyoon Choi. Overview of informetrics research activities in KISTI Knowledge Information Center. Collnet Journal of Scientometrics and Information Management, 6(1):v, 2012. CODEN ???? ISSN 0973-7766 (print), 2168-930X (electronic).

Choi:2013:KSI

[CKJC13] Honam Choi, Byungkyu Kim, Youngim Jung, and Seonheui Choi. Korean scholarly information analysis based on Korea Science Citation Database (KSCD). Collnet Journal of Scientometrics and Information Management, 7(1):1–33, 2013. CODEN ???? ISSN 0973-7766 (print), 2168-930X (electronic).

Chatterjee:2016:FFI

[CKM16] Usashi Chatterjee, Vinit Kumar, and Devika P. Madalli. Formalizing food ingredients for data analysis and knowledge organization. Collnet Journal of Scientometrics and Information Management, 10(2):289–309, 2016. CODEN ???? ISSN 0973-7766 (print), 2168-930X (electronic). URL http:

//www.tandfonline.com/doi/abs/10.1080/09737766.2016.
1213970.

Ma:2014:QCR

[cMhLyW14] Fei cheng Ma, Peng hui Lyu, and Xiao yong Wang. Quantitative characterization of the research productivity of interdisciplinary. Collnet Journal of Scientometrics and Information Management, 8(1):153–168, 2014. CODEN ???? ISSN 0973-7766 (print), 2168-930X (electronic).

Lo:2014:REC

[cSL14] Szu chia S. Lo. Re-examination contribution of academic research to development of industrial technologies. Collnet Journal of Scientometrics and Information Management, 8(1):123–140, 2014. CODEN ???? ISSN 0973-7766 (print), 2168-930X (electronic).

Chhatoi:2020:KMD

[CSN20] Biswajit Prasad Chhatoi, Sharada Prasad Sahoo, and Durga Prasad Nayak. Knowledge mapping in the discipline of financial literacy and charting the future directions. Collnet Journal of Scientometrics and Information Management, 14(2):311–330, 2020. CODEN ???? ISSN 0973-7766 (print), 2168-930X

(electronic). URL http://www.tandfonline.com/doi/abs/10.
1080/09737766.2020.1841584.

Chen:2011:BAS

[CWLL11] Yue Chen, Xianwen Wang, Deming Lin, and Zeyuan Liu. A bibliometric analysis of science cooperation group size. Collnet Journal of Scientometrics and Information Management, 5(2): 227–237, 2011. CODEN ???? ISSN 0973-7766 (print), 2168-930X (electronic).

Chen:2020:ALZ

[CWS20] Yue Chen, Zhiqi Wang, and Kai Song. The academic life of Zeyuan Liu. Collnet Journal of Scientometrics and Information Management, 14(1):3–7, 2020. CODEN ???? ISSN 0973-7766 (print), 2168-930X (electronic). URL http:

//www.tandfonline.com/doi/abs/10.1080/09737766.2020.
1830730.

Zhou:2020:ETM

[CZ20] Jin Chen and Wei Zhou. The exploration of travel motivation research: a scientometric analysis based on CiteSpace. Collnet Journal of Scientometrics and Information Management, 14(2): 257–283, 2020. CODEN ???? ISSN 0973-7766 (print), 2168-930X

(electronic). URL http://www.tandfonline.com/doi/abs/10.
1080/09737766.2021.1906779.

Darvish:2012:ADN

[Dar12] Hamid Darvish. Assessing the diffusion of nanotechnology in Turkey: A social network analysis approach. Collnet Journal of Scientometrics and Information Management, 6(1):175–183, 2012. CODEN ???? ISSN 0973-7766 (print), 2168-930X (electronic).

Beaver:2008:OR

[dB08] Donald deB. Beaver. Opening remarks. Collnet Journal of Scientometrics and Information Management, 2(1):iii, 2008. CODEN ???? ISSN 0973-7766 (print), 2168-930X (electronic).

Beaver:2013:MFC

[dB13] Donald deB. Beaver. The many faces of collaboration and teamwork in scientific research: Updated reflections on scientific collaboration. Collnet Journal of Scientometrics and Information Management, 7(1):45–54, 2013. CODEN ???? ISSN 0973-7766 (print), 2168-930X (electronic).

Filho:2010:RSA

[dCPFA10] Roberto de Camargo Penteado Filho and Antonio Flavio Dias Avila. Reflections on a scientometrical analysis of Embrapa Brazil Web of Science articles from 1977 to 2006. Collnet Journal of Scientometrics and Information Management, 4(1):9–20, 2010. CODEN ???? ISSN 0973-7766 (print), 2168-930X (electronic).

Dhawan:2018:SSA

[DGG18] S. M. Dhawan, B. M. Gupta, and Ritu Gupta. Supercomputing: a scientometric assessment of global publications output during 2007–16. Collnet Journal of Scientometrics and Information Management, 12(2):197–213, 2018. CODEN ???? ISSN 0973-7766 (print), 2168-930X (electronic). URL http:

//www.tandfonline.com/doi/abs/10.1080/09737766.2018.
1545395.

Deming:2009:ICK

[DHZK09] Lin Deming, Hou Haiyan, Liu Zeyuan, and Ding Kun. The International Conference of Knowledge Visualization — summary of the 5^{th} International Conference on Webometrics, Informetrics and Scientometrics & 10^{th} COLLNET Meeting. Collnet Journal of Scientometrics and Information Management, 3(2):1–2, 2009. CODEN ???? ISSN 0973-7766 (print), 2168-930X (electronic).

Dissanayake:2015:MPR

[Dis15] H. N. K. Dissanayake. Medicinal plant research in Sri Lanka: A scientometric study based on Scopus database. Collnet Journal of Scientometrics and Information Management, 9(2):225– 234, 2015. CODEN ???? ISSN 0973-7766 (print), 2168-930X

(electronic). URL http://www.tandfonline.com/doi/abs/10.
1080/09737766.2015.1069961.

Santos:2021:SMS

[dJSdJP21] Jânio Rodrigo de Jesus Santos, Carlos Augusto Francisco de Jesus, and Cláudio Damasceno Pinto. Scientific mapping of stem cells associated with Chagas disease: a bibliometric analysis. Collnet Journal of Scientometrics and Information Management, 15(2):429–443, 2021. CODEN ???? ISSN 0973-7766 (print),

2168-930X (electronic). URL http://www.tandfonline.com/
doi/abs/10.1080/09737766.2021.1977094.

Santos:2021:TSU

[dJSPR21] Jânio Rodrigo de Jesus Santos, Cláudio Damasceno Pinto, and Angela Machado Rocha. Technological scenarios of the use of nanobiotechnology in strategies against Zika virus. Collnet Journal of Scientometrics and Information Management, 15(2):413– 428, 2021. CODEN ???? ISSN 0973-7766 (print), 2168-930X

(electronic). URL http://www.tandfonline.com/doi/abs/10.
1080/09737766.2021.1981172.

Dora:2017:EAR

[DK17] Mallikarjun Dora and H. Anil Kumar. An empirical analysis of the research trends in the field of library and information science in India — 2004–2015. Collnet Journal of Scientometrics and Information Management, 11(2):361–378, 2017. CODEN ???? ISSN 0973-7766 (print), 2168-930X

(electronic). URL http://www.tandfonline.com//doi/abs/
10.1080/09737766.2017.1317959.

Dutt:2010:SPG

[DKG10] Bharvi Dutt, Suresh Kumar, and K. C. Garg. Scientometric profile of global dengue research. Collnet Journal of Scientometrics and Information Management, 4(1):81–91, 2010. CODEN ???? ISSN 0973-7766 (print), 2168-930X (electronic).

Dhoble:2018:PPI

[DKK18] Shilpa Dhoble, S. Kumar, and Surendra Kumar. Publication productivity of Indian scientists in groundnut research: A bibliometric study. Collnet Journal of Scientometrics and Information Management, 12(1):149–159, 2018. CODEN ???? ISSN 0973-7766 (print), 2168-930X (electronic). URL http:

//www.tandfonline.com//doi/abs/10.1080/09737766.2018.
1433090.

Daud:2014:CAC

[DM14] Ali Daud and Faqir Muhammad. Consistent annual citations based researcher index. Collnet Journal of Scientometrics and Information Management, 8(2):209–216, 2014. CODEN ???? ISSN 0973-7766 (print), 2168-930X (electronic).

Darmadji:2018:PIU

[DPR^{+}18] Ahmad Darmadji, Lantip Diat Prasojo, Yatim Riyanto, Fitri Ayu Kusumaningrum, and Yuli Andriansyah. Publications of Islamic University of Indonesia in Scopus database: A bibliometric assessment. Collnet Journal of Scientometrics and Information Management, 12(1):109–131, 2018. CODEN ???? ISSN 0973-7766 (print), 2168-930X (electronic). URL http:

//www.tandfonline.com//doi/abs/10.1080/09737766.2017.
1400754.

Dubey:2020:QAN

[DSM20] Shubham Dubey, Prasun Sharma, and Borbély Mária. Quality assurance of national digital library of University of Debrecen: an extensive association rules based study using ISO/IEC 25010. Collnet Journal of Scientometrics and Information Management, 14(1):119–133, 2020. CODEN ???? ISSN 0973-7766 (print),

2168-930X (electronic). URL http://www.tandfonline.com/
doi/abs/10.1080/09737766.2020.1819172.

Vanz:2012:SOI

[dSVS12] Samile Andréa de Souza Vanz and Ida Regina Chittó Stumpf. Scientific output indicators and scientific collaboration network mapping in Brazil. Collnet Journal of Scientometrics and Information Management, 6(2):315–334, 2012. CODEN ???? ISSN 0973-7766 (print), 2168-930X (electronic).

Devi:2018:WCD

[DV18] Ksh. Krishna Devi and Manoj Kumar Verma. Web content and design trends of Indian Institute of Technology (IITs) libraries’ website: An evaluation. Collnet Journal of Scientometrics and Information Management, 12(2):165–181, 2018. CODEN ???? ISSN 0973-7766 (print), 2168-930X

(electronic). URL http://www.tandfonline.com/doi/abs/10.
1080/09737766.2018.1433100.

Elango:2016:SCN

[EBS16] B. Elango, Lutz Bornmann, and S. Shankar. Study of citation networks in tribology research. Collnet Journal of Scientometrics and Information Management, 10(1):71–96, 2016. CODEN ???? ISSN 0973-7766 (print), 2168-930X (electronic).

Emebo:2020:IIA

[EE20] Azubuike Ezenwoke and Onyeka Emebo. The impact of internet access on cloud computing research in Africa: Analysis of bibliometric and online search data. Collnet Journal of Scientometrics and Information Management, 14(2):393–410, 2020. CODEN ???? ISSN 0973-7766 (print), 2168-930X

(electronic). URL http://www.tandfonline.com/doi/abs/10.
1080/09737766.2021.1927881.

Egghe:2013:DAI

[Egg13] L. Egghe. Dynamical aspects of information production processes. Collnet Journal of Scientometrics and Information Management, 7(1):35–44, 2013. CODEN ???? ISSN 0973-7766 (print), 2168-930X (electronic).

Erfanmanesh:2013:PIP

[EGJ13] Mohammadamin Erfanmanesh, Zohreh Gholamhosseinzadeh, and Reza Basirian Jahromi. The performance of Iranian psychology and psychiatry researchers: A scientometric study. Collnet Journal of Scientometrics and Information Management, 7 (2):161–172, 2013. CODEN ???? ISSN 0973-7766 (print), 2168-930X (electronic).

Egghe:2008:CPI

[EGK08] Leo Egghe, Marc Goovaerts, and Hildrun Kretschmer. Collaboration and productivity: an investigation into ‘Scientometrics’ journal and ‘UHasselt’ repository. Collnet Journal of Scientometrics and Information Management, 2(1):83–89, 2008. CODEN ???? ISSN 0973-7766 (print), 2168-930X (electronic).

Elango:2018:TCA

[EH18] Bakthavachalam Elango and Yuh Shan Ho. Top-cited articles in the field of tribology: a bibliometric analysis. Collnet Journal of Scientometrics and Information Management, 12(2):289– 307, 2018. CODEN ???? ISSN 0973-7766 (print), 2168-930X

(electronic). URL http://www.tandfonline.com/doi/abs/10.
1080/09737766.2018.1529125.

Erfanmanesh:2013:SPI

[EJHG13] Mohammadamin Erfanmanesh, Reza Basirian Jahromi, Elahe Hosseini, and Zohreh Gholamhosseinzadeh. Scientific productivity, impact and collaboration of the top Asian countries in Scopus during 1996–2010. Collnet Journal of Scientometrics and Information Management, 7(1):97–110, 2013. CODEN ???? ISSN 0973-7766 (print), 2168-930X (electronic).

Emami:2018:TSP

[EKM^{+}18] Zahra Emami, Mohammad Ebrahim Khamseh, Nahid Hashemi Madani, Nadjla Hariri, Mohammad Reza Alibeyk, and Omolbanin Asadi Ghadiqolaei. Trend of scientific productions in the field of endocrinology and metabolism in Middle East countries during 2007–2013. Collnet Journal of Scientometrics and Information Management, 12(1):55–71, 2018. CODEN ???? ISSN 0973-7766 (print), 2168-930X (electronic). URL http:

//www.tandfonline.com//doi/abs/10.1080/09737766.2017.
1400751.

Elango:2019:CEB

[Ela19] Bakthavachalam Elango. Calculation errors in bibliometrics: The case of CAGR. Collnet Journal of Scientometrics and Information Management, 13(2):331–337, 2019. CODEN ???? ISSN 0973-7766 (print), 2168-930X (electronic). URL http:

//www.tandfonline.com/doi/abs/10.1080/09737766.2019.
1694397.

Elango:2020:NIC

[Ela20] Bakthavachalam Elango. A new indicator for comparing the rankings: The Relative Rank Index (RRI). Collnet Journal of Scientometrics and Information Management, 14(1):153–161, 2020. CODEN ???? ISSN 0973-7766 (print), 2168-930X

(electronic). URL http://www.tandfonline.com/doi/abs/10.
1080/09737766.2020.1816512.

Ebadi:2015:ASA

[ES15] Ashkan Ebadi and Andrea Schiffauerova. Analyzing scientific activities of the top ten Canadian universities. Collnet Journal of Scientometrics and Information Management, 9(1):93–106, 2015. CODEN ???? ISSN 0973-7766 (print), 2168-930X (electronic).

Elshawi:2016:ICC

[ES16a] Radwa Elshawi and Sherif Sakr. International conferences on computer system: Analysis of EuroSys, SOSP, and OSDI during 2006–2014. Collnet Journal of Scientometrics and Information Management, 10(1):175–195, 2016. CODEN ???? ISSN 0973-7766 (print), 2168-930X (electronic).

Elshawi:2016:AIA

[ES16b] Radwa Elshawi and Sherif Sakr. On analyzing the impact of authors and their collaboration patterns in the major computer algorithms research conferences. Collnet Journal of Scientometrics and Information Management, 10(1):155–173, 2016. CO-DEN ???? ISSN 0973-7766 (print), 2168-930X (electronic).

Soboleva:2020:OCR

[ES20] Yulia B. Evdokimenkova and Natalya O. Soboleva. Organic chemistry in Russia: Bibliometric publication flow analysis over the past 30 years. Collnet Journal of Scientometrics and Information Management, 14(1):23–36, 2020. CODEN ???? ISSN 0973-7766 (print), 2168-930X (electronic). URL http:

//www.tandfonline.com/doi/abs/10.1080/09737766.2020.
1766388.

Fakruddin:2014:JIF

[Fak14] Md. Fakruddin. Journal Impact Factor — applicability and alternative indices. Collnet Journal of Scientometrics and Information Management, 8(2):457–464, 2014. CODEN ???? ISSN 0973-7766 (print), 2168-930X (electronic).

Fanourgiakis:2021:ERB

[FPK21] J. Fanourgiakis, C. Panagiotakis, and I. Konstantourakis. Exploring the research-based ranking and the university impact in Greece: a case of peripheral departments of management and economics. Collnet Journal of Scientometrics and Information Management, 15(2):305–322, 2021. CODEN ???? ISSN 0973-7766 (print), 2168-930X (electronic). URL http:

//www.tandfonline.com/doi/abs/10.1080/09737766.2021.
1978826.

Garfield:2012:CCI

[Gar12] Eugene Garfield. A century of citation indexing. Collnet Journal of Scientometrics and Information Management, 6(1):1–6, 2012. CODEN ???? ISSN 0973-7766 (print), 2168-930X (electronic).

Gupta:2011:MTR

[GB11] B. M. Gupta and Adarsh Bala. Mapping of tuberculosis research in India: A scientometric analysis of publications output during 1998–2009. Collnet Journal of Scientometrics and Information Management, 5(1):33–51, 2011. CODEN ???? ISSN 0973-7766 (print), 2168-930X (electronic).

Bebi:2021:CJS

[GB21] K. C. Garg and Bebi. COLLNET Journal of Scientometrics and information Management: a bibliometric study. Collnet Journal of Scientometrics and Information Management, 15(1):47– 61, 2021. CODEN ???? ISSN 0973-7766 (print), 2168-930X

(electronic). URL http://www.tandfonline.com/doi/abs/10.
1080/09737766.2021.1920067.

Gupta:2013:WRO

[GBG13] B. M. Gupta, Adarsh Bala, and Ritu Gupta. World research output in conjunctivitis research: A quantitative analysis during 2002–11. Collnet Journal of Scientometrics and Information Management, 7(2):261–275, 2013. CODEN ???? ISSN 0973-7766 (print), 2168-930X (electronic).

Gupta:2011:MAH

[GBK11] B. M. Gupta, Adarsh Bala, and Har Kaur. Mapping of AIDS/ HIV research in India: A scientometrics analysis of publications output during 1999–2008. Collnet Journal of Scientometrics and Information Management, 5(2):185–203, 2011. CODEN ???? ISSN 0973-7766 (print), 2168-930X (electronic).

Gupta:2019:GRS

[GD19a] B. M. Gupta and S. M. Dhawan. Global research studies on “electronic resources in libraries”: a scientometric assessment during 1994–2017. Collnet Journal of Scientometrics and Information Management, 13(1):167–181, 2019. CODEN ???? ISSN 0973-7766 (print), 2168-930X (electronic). URL http:

//www.tandfonline.com/doi/abs/10.1080/09737766.2019.
1583836.

Gupta:2019:PCR

[GD19b] B. M. Gupta and S. M. Dhawan. Parallel computing research in India: a scientometric assessment of Indian publications output during 1998–2017. Collnet Journal of Scientometrics and Information Management, 13(1):111–126, 2019. CODEN ???? ISSN 0973-7766 (print), 2168-930X (elec-

tronic). URL http://www.tandfonline.com/doi/abs/10.
1080/09737766.2018.1560638.

Gu:2014:MPS

[GDY14] Li Gu, Kun Ding, and Weichun Yan. Management or productivity: A scientometric analysis on the role of presidents of 6 universities. Collnet Journal of Scientometrics and Information Management, 8(1):81–90, 2014. CODEN ???? ISSN 0973-7766 (print), 2168-930X (electronic).

Gupta:2014:ILR

[GG14] Ritu Gupta and B. M. Gupta. Indian lymphoma research: A scientometric analysis of Indian publications output during 2004– 13. Collnet Journal of Scientometrics and Information Management, 8(2):419–436, 2014. CODEN ???? ISSN 0973-7766 (print), 2168-930X (electronic).

Garg:2016:SPI

[GK16] K. C. Garg and S. Kumar. Scientometric profile of an Indian state: The case study of Odisha. Collnet Journal of Scientometrics and Information Management, 10(1):141–153, 2016. CO-DEN ???? ISSN 0973-7766 (print), 2168-930X (electronic).

Garg:2019:CIW

[GK19] K. C. Garg and S. Kumar. Contribution of Indian women scholars to Indian output in bibliometrics/scientometrics and the pattern of co-authorship and collaboration. Collnet Journal of Scientometrics and Information Management, 13(1):127– 138, 2019. CODEN ???? ISSN 0973-7766 (print), 2168-930X

(electronic). URL http://www.tandfonline.com/doi/abs/10.
1080/09737766.2018.1564590.

Guo:2008:DCA

[GKL08] Hanning Guo, Hildrun Kretschmer, and Zeyuan Liu. Distribution of co-author pairs’ frequencies of the Journal of Information Technology. Collnet Journal of Scientometrics and Information Management, 2(1):73–81, 2008. CODEN ???? ISSN 0973-7766 (print), 2168-930X (electronic).

Glanzel:2008:SMB

[Glä08] Wolfgang Glänzel. Seven myths in bibliometrics about facts and fiction in quantitative science studies. Collnet Journal of Scientometrics and Information Management, 2(1):9–17, 2008. CODEN ???? ISSN 0973-7766 (print), 2168-930X (electronic).

Glanzel:2011:BCH

[Glä11] Wolfgang Glänzel. A brief communication: How to obtain Erdős number 1 in the 21^{st} Century? Some thoughts on posthumous coauthorship. Collnet Journal of Scientometrics and Information Management, 5(1):1–4, 2011. CODEN ???? ISSN 0973-7766 (print), 2168-930X (electronic).

Garfield:2015:JIF

[GP15] Eugene Garfield and Alexander Pudovkin. Journal Impact Factor strongly correlates with the citedness of the median journal paper. Collnet Journal of Scientometrics and Information Management, 9(1):5–14, 2015. CODEN ???? ISSN 0973-7766 (print), 2168-930X (electronic).

Roy:2021:WAO

[GR21] Sanjib Ghosh and Bijan Kumar Roy. Webometric analysis of open access digital repository of agricultural sciences in Asia. Collnet Journal of Scientometrics and Information Management, 15(1):107–132, 2021. CODEN ???? ISSN 0973-7766 (print),

2168-930X (electronic). URL http://www.tandfonline.com/
doi/abs/10.1080/09737766.2021.1938743.

Gu:2015:WAM

[GRAY15] Li Gu, Liqiang Ren, Shule An, and Weichun Yan. The webometrics analysis of market share and network attention of operation systems for personal computers — based on the Baidu index as a Web metrics research. Collnet Journal of Scientometrics and Information Management, 9(2):161–173, 2015. CODEN ???? ISSN 0973-7766 (print), 2168-930X

(electronic). URL http://www.tandfonline.com/doi/abs/10.
1080/09737766.2015.1069953.

Gupta:2010:SSK

[Gup10] B. M. Gupta. Status of South Korea in science and technology as reflected in its publication output, 1997–2007. Collnet Journal of Scientometrics and Information Management, 4(1):53–79, 2010. CODEN ???? ISSN 0973-7766 (print), 2168-930X (electronic).

Wani:2021:BAV

[GW21] Shabir Ahmad Ganaie and Javaid Ahmad Wani. Bibliometric analysis and visualization of Nanotechnology Research Field. Collnet Journal of Scientometrics and Information Management, 15(2):445–467, 2021. CODEN ???? ISSN 0973-7766 (print),

2168-930X (electronic). URL http://www.tandfonline.com/
doi/abs/10.1080/09737766.2021.2008780.

Hassanzadeh:2015:RBD

[HA15] Mohammad Hassanzadeh and Babak Akhgar. Relationship between development indicators and contribution to the science: Experiences from Iran. Collnet Journal of Scientometrics and Information Management, 9(2):311–319, 2015. CODEN ???? ISSN 0973-7766 (print), 2168-930X (elec-

tronic). URL http://www.tandfonline.com/doi/abs/10.
1080/09737766.2015.1069969.

Hassan:2021:BAJ

[HAM^{+}21] M. Kabir Hassan, Muneer M. Alshater, Jinnatul Raihan Mumu, Adel M. Sarea, and Md. Abul Kalam Azad. Bibliometric analysis of the Journal of Islamic Accounting and Business Research: Ten years review. Collnet Journal of Scientometrics and Information Management, 15(1):63–88, 2021. CODEN ???? ISSN 0973-7766 (print), 2168-930X (elec-

tronic). URL http://www.tandfonline.com/doi/abs/10.
1080/09737766.2021.1934603.

Hassanzadeh:2014:IKS

[HAN14] Mohammad Hassanzadeh, Babak Akhgar, and Fatemeh Navidi. Integrating knowledge and science indicators (Kientometrics) towards a rational framework of investigations: A comparative approach. Collnet Journal of Scientometrics and Information Management, 8(2):247–256, 2014. CODEN ???? ISSN 0973-7766 (print), 2168-930X (electronic).

Harnisch:2014:IDD

[Har14] Michael J. Harnisch. Impacts of the developments in digital media and communication technology on personalization research. Collnet Journal of Scientometrics and Information Management, 8(2):353–370, 2014. CODEN ???? ISSN 0973-7766 (print), 2168-930X (electronic).

Batcha:2021:SAR

[HB21] Saddam Hossain and M. Sadik Batcha. Scientometric analysis of research productivity from Indian dialysis over the last twenty years in Web of Science. Collnet Journal of Scientometrics and Information Management, 15(2):323–339, 2021. CODEN ???? ISSN 0973-7766 (print), 2168-930X

(electronic). URL http://www.tandfonline.com/doi/abs/10.
1080/09737766.2021.2005455.

Herubel:2008:FDP

[Hér08] Jean-Pierre V. M. Hérubel. French dissertations in a period of transition, 1971–1990: Historical and general observations for information and communication sciences (SIC), philosophy, and art and archaeology. Collnet Journal of Scientometrics and Information Management, 2(2):71–77, 2008. CODEN ???? ISSN 0973-7766 (print), 2168-930X (electronic).

Herubel:2012:SSS

[Hér12] Jean-Pierre V. M. Hérubel. Subject specialization and subject dispersion in historical scholarship: Qualitative preliminary explorations in disciplinary cultures. Collnet Journal of Scientometrics and Information Management, 6(1):131–144, 2012. CO-DEN ???? ISSN 0973-7766 (print), 2168-930X (electronic).

Hartley:2020:BAH

[HH20] Yuh-Shan Ho and James Hartley. A bibliometric analysis of highly cited Phosphoinositide 3-Kinase (PI3K) research papers. Collnet Journal of Scientometrics and Information Management, 14(1):37–54, 2020. CODEN ???? ISSN 0973-7766 (print), 2168-

930X (electronic). URL http://www.tandfonline.com/doi/
abs/10.1080/09737766.2020.1768172.

Jeong:2013:ATC

[hJjJiK13] Dae hyun Jeong, Hye jeong Jang, and Young il Kwon. Analysis on technical competitiveness among major applicants in secondary battery fields. Collnet Journal of Scientometrics and Information Management, 7(1):81–96, 2013. CODEN ???? ISSN 0973-7766 (print), 2168-930X (electronic).

Husain:2008:EDa

[HK08a] Shabahat Husain and Hildrun Kretschmer. From Editors’ desk. Collnet Journal of Scientometrics and Information Management, 2(1):i, 2008. CODEN ???? ISSN 0973-7766 (print), 2168-930X (electronic).

Husain:2008:EDb

[HK08b] Shabahat Husain and Hildrun Kretschmer. From Editors’ desk. Collnet Journal of Scientometrics and Information Management, 2(2):i, 2008. CODEN ???? ISSN 0973-7766 (print), 2168-930X (electronic).

Husain:2009:EDa

[HK09a] Shabahat Husain and Hildrun Kretschmer. From Editors’ desk. Collnet Journal of Scientometrics and Information Management, 3(1):i, 2009. CODEN ???? ISSN 0973-7766 (print), 2168-930X (electronic).

Husain:2009:EDb

[HK09b] Shabahat Husain and Hildrun Kretschmer. From Editors’ desk. Collnet Journal of Scientometrics and Information Management, 3(2):i, 2009. CODEN ???? ISSN 0973-7766 (print), 2168-930X (electronic).

Husain:2010:EDa

[HK10] Shabahat Husain and Hildrun Kretschmer. From Editors’ desk. Collnet Journal of Scientometrics and Information Management, 4(1):i, 2010. CODEN ???? ISSN 0973-7766 (print), 2168-930X (electronic).

Hou:2009:ICN

[HLK^{+}09] Haiyan Hou, Zeyuan Liu, Hildrun Kretschmer, Tianpeng Qu, and Chunting Lu. International collaboration networks of Chinese scientometrics. Collnet Journal of Scientometrics and Information Management, 3(1):61–70, 2009. CODEN ???? ISSN 0973-7766 (print), 2168-930X (electronic).

Husain:2010:ARI

[HM10] Shabahat Husain and Muzamil Mushtaq. Academic ranking of Indian central universities in science and technology as represented in Science Citation Index — Expanded: Scientometric study (2005–2009). Collnet Journal of Scientometrics and Information Management, 4(2):81–90, 2010. CODEN ???? ISSN 0973-7766 (print), 2168-930X (electronic).

Hassanzadeh:2009:OSI

[HNN09] M. Hassanzadeh, G. Nourmohammadi, and A. Noroozichakoli. Organization of science in Iran: A holistic review. Collnet Journal of Scientometrics and Information Management, 3(2):53–59, 2009. CODEN ???? ISSN 0973-7766 (print), 2168-930X (electronic).

Ho:2018:CPP

[Ho18] Yuh-Shan Ho. Classic papers published by Taiwanese scientists in the science citation index expanded: A bibliometric study. Collnet Journal of Scientometrics and Information Management, 12(1):83–95, 2018. CODEN ???? ISSN 0973-7766 (print), 2168-

930X (electronic). URL http://www.tandfonline.com//doi/
abs/10.1080/09737766.2017.1400752.

Holmberg:2011:LBW

[Hol11] Kim Holmberg. Linking to a bilingual Web space. Collnet Journal of Scientometrics and Information Management, 5(1):5–16, 2011. CODEN ???? ISSN 0973-7766 (print), 2168-930X (electronic).

Husain:2010:EDb

[HRK10] Shabahat Husain, I. K. Ravichandra Rao, and Hildrun Kretschmer. From Editors’ desk. Collnet Journal of Scientometrics and Information Management, 4(2):i, 2010. CODEN ???? ISSN 0973-7766 (print), 2168-930X (electronic).

Hasan:2015:LIS

[HS15] Nabi Hasan and Mukhtiar Singh. Library and information science research output: A study based on Web of Science. Collnet Journal of Scientometrics and Information Management, 9(1): 47–64, 2015. CODEN ???? ISSN 0973-7766 (print), 2168-930X (electronic).

Husain:2007:EDa

[Hus07a] Shabahat Husain. From Editor’s desk. Collnet Journal of Scientometrics and Information Management, 1(1):iii, 2007. CODEN ???? ISSN 0973-7766 (print), 2168-930X (electronic).

Husain:2007:EDb

[Hus07b] Shabahat Husain. From Editor’s desk. Collnet Journal of Scientometrics and Information Management, 1(2):iii, 2007. CODEN ???? ISSN 0973-7766 (print), 2168-930X (electronic).

Husain:2009:BSP

[Hus09] Shabahat Husain. Biographical sketch of Prof. Shabahat Husain. Collnet Journal of Scientometrics and Information Management, 3(2):v, 2009. CODEN ???? ISSN 0973-7766 (print), 2168-930X (electronic).

Wang:2020:BAA

[HW20] Yuh-Shan Ho and Ming-Huang Wang. A bibliometric analysis of artificial intelligence publications from 1991 to 2018. Collnet Journal of Scientometrics and Information Management, 14(2): 369–392, 2020. CODEN ???? ISSN 0973-7766 (print), 2168-930X

(electronic). URL http://www.tandfonline.com/doi/abs/10.
1080/09737766.2021.1918032.

Hou:2013:DSC

[HWL^{+}13] Haiyan Hou, Chuanli Wang, Chunjuan Luan, Xianwen Wang, and Ping Zhuang. The dynamics of scientific collaboration networks in scientometrics. Collnet Journal of Scientometrics and Information Management, 7(1):121–140, 2013. CODEN ???? ISSN 0973-7766 (print), 2168-930X (electronic).

Hsu:2020:BAH

[HYWH20] Y-Hsin Elsa Hsu, Ya-Ting Yang, Ming-Huang Wang, and Yuh-Shan Ho. Bibliometric analysis of highly cited publications in health policy and services. Collnet Journal of Scientometrics and Information Management, 14(1):177–202, 2020. CODEN ???? ISSN 0973-7766 (print), 2168-930X

(electronic). URL http://www.tandfonline.com/doi/abs/10.
1080/09737766.2020.1816513.

Kwon:2014:TRA

[iKhJ14] Young il Kwon and Dae hyun Jeong. Technology relevance analysis between wind power energy — fuel cell — green car using network analysis, IPC map. Collnet Journal of Scientometrics and Information Management, 8(1):109–121, 2014. CODEN ???? ISSN 0973-7766 (print), 2168-930X (electronic).

Ivancheva:2008:STM

[Iva08] Ludmila Ivancheva. Scientometrics today: A methodological overview. Collnet Journal of Scientometrics and Information Management, 2(2):47–56, 2008. CODEN ???? ISSN 0973-7766 (print), 2168-930X (electronic).

Jahangiri:2014:ISA

[JAAN14] Saeideh Jahangiri, Amir Reza Asnafi, Amir Hussein Rajabzadeh Assarha, and Maryam Pakdaman Naeini. Iranian scientific associations’ websites in the field of humanities: A Webometric study. Collnet Journal of Scientometrics and Information Management, 8(2):273–279, 2014. CODEN ???? ISSN 0973-7766 (print), 2168-930X (electronic).

Jalal:2010:WPS

[JBM10] Samir Kumar Jalal, Subal Chandra Biswas, and Parthasarathi Mukhopadhyay. Web presence of selected Asian countries: A Webometric study. Collnet Journal of Scientometrics and Information Management, 4(2):57–68, 2010. CODEN ???? ISSN 0973-7766 (print), 2168-930X (electronic).

Jung:2013:SJU

[JKK13] Youngim Jung, Jayhoon Kim, and Hwanmin Kim. STM e-journal use analysis by utilizing KESLI usage statistics consolidation platform. Collnet Journal of Scientometrics and Information Management, 7(2):205–215, 2013. CODEN ???? ISSN 0973-7766 (print), 2168-930X (electronic).

Jallow:2021:CAA

[JLH21] Amadou W. Jallow, Shin-Da Lee, and Yuh-Shan Ho. Classic articles in apoptotic research: a bibliometric analysis. Collnet Journal of Scientometrics and Information Management, 15(1): 145–162, 2021. CODEN ???? ISSN 0973-7766 (print), 2168-930X

(electronic). URL http://www.tandfonline.com/doi/abs/10.
1080/09737766.2021.1934605.

Jhamb:2019:IGR

[JMS19] Garvita Jhamb, Meera, and K. P. Singh. Indian geology research as reflected by Web of Science during 1998–2017. Collnet Journal of Scientometrics and Information Management, 13(1): 37–51, 2019. CODEN ???? ISSN 0973-7766 (print), 2168-930X

(electronic). URL http://www.tandfonline.com/doi/abs/10.
1080/09737766.2019.1568363.

Janmaijaya:2020:EER

[JSM20] Manvendra Janmaijaya, Amit K. Shukla, and Pranab K. Muhuri. An evaluation of the economics research in India during last three decades (1989-2019). Collnet Journal of Scientometrics and Information Management, 14(1):93–118, 2020. CODEN ???? ISSN 0973-7766 (print), 2168-930X

(electronic). URL http://www.tandfonline.com/doi/abs/10.
1080/09737766.2020.1802208.

Jallow:2020:GRT

[JWH20] Amadou W. Jallow, Ming-Huang Wang, and Yuh-Shan Ho. Global research trends and publications of insomnia: a bibliometric analysis. Collnet Journal of Scientometrics and Information Management, 14(2):349–367, 2020. CODEN ???? ISSN 0973-7766 (print), 2168-930X (electronic). URL http:

//www.tandfonline.com/doi/abs/10.1080/09737766.2021.
1906184.

Krishnamoorthy:2008:HSI

[KA08] G. Krishnamoorthy and A. Amudhavalli. Health sciences in India: a scientometric study. Collnet Journal of Scientometrics and Information Management, 2(2):79–85, 2008. CODEN ???? ISSN 0973-7766 (print), 2168-930X (electronic).

Kalita:2016:CAS

[Kal16] Deepjyoti Kalita. Citation analysis of Science. Collnet Journal of Scientometrics and Information Management, 10(2):237– 254, 2016. CODEN ???? ISSN 0973-7766 (print), 2168-930X

(electronic). URL http://www.tandfonline.com/doi/abs/10.
1080/09737766.2016.1213967.

Kundra:2008:SCA

[KdBKK08] Ramesh Kundra, Donald B. de Beaver, Hildrun Kretschmer, and Theo Kretschmer. Studies in co-authorship pairs distribution: Part 2. Co-author pairs’ frequencies distribution in journals of gender studies. Collnet Journal of Scientometrics and Information Management, 2(1):63–71, 2008. CODEN ???? ISSN 0973-7766 (print), 2168-930X (electronic).

Kaur:2009:ICP

[KG09] Har Kaur and B. M. Gupta. Indian contribution in pharmacology, toxicology & pharmaceutics during 1998–2007: A scientometric analysis. Collnet Journal of Scientometrics and Information Management, 3(1):1–9, 2009. CODEN ???? ISSN 0973-7766 (print), 2168-930X (electronic).

Karpagam:2011:PTN

[KGB11] R. Karpagam, S. Gopalakrishnan, and B. Ramesh Babu. Publication trend on Nanotechnology among G15 countries: A bibliometric study. Collnet Journal of Scientometrics and Information Management, 5(1):61–80, 2011. CODEN ???? ISSN 0973-7766 (print), 2168-930X (electronic).

Karpagam:2012:SAS

[KGBN12] R. Karpagam, S. Gopalakrishnan, B. Ramesh Babu, and M. Natarajan. Scientometric analysis of stem cell research: A comparative study of India and other countries. Collnet Journal of Scientometrics and Information Management, 6(2):229–252, 2012. CODEN ???? ISSN 0973-7766 (print), 2168-930X (electronic).

Kwon:2013:APA

[KhJ13] Young-Il Kwon and Dae hyun Jeong. Analysis on patent activity and technical diversity in green car fields. Collnet Journal of Scientometrics and Information Management, 7(1):141–159, 2013. CODEN ???? ISSN 0973-7766 (print), 2168-930X (electronic).

Kretschmer:2016:RIC

[KJ16] Hildrun Kretschmer and P. K. Jain. Report of The 11th International Conference on Webometrics, Informetrics and Scientometrics (WIS) & 16th COLLNET Meeting November 26-28, 2015, New Delhi, India [Hosted by Institute of Economic Growth, Uni-

versity of Delhi, India] http://www.slp.org.in/collnet2015.

Collnet Journal of Scientometrics and Information Management, 10(1):1–7, 2016. CODEN ???? ISSN 0973-7766 (print), 2168-930X (electronic).

Kretschmer:2007:DCA

[KK07a] Hildrun Kretschmer and Theo Kretschmer. Distribution of coauthor pairs’ frequencies of the journal of biological chemistry explained as social gestalt. Collnet Journal of Scientometrics and Information Management, 1(2):21–25, 2007. CODEN ???? ISSN 0973-7766 (print), 2168-930X (electronic).

Kretschmer:2007:NCM

[KK07b] Hildrun Kretschmer and Theo Kretschmer. A new centrality measure for social network analysis applicable to bibliometric and Webometric data. Collnet Journal of Scientometrics and Information Management, 1(1):1–7, 2007. CODEN ???? ISSN 0973-7766 (print), 2168-930X (electronic).

Kretschmer:2008:SCA

[KK08] Hildrun Kretschmer and Theo Kretschmer. Studies in coauthorship pairs distribution: Part 1. Varying shapes of coauthor pairs’ distributions. Collnet Journal of Scientometrics and Information Management, 2(1):45–61, 2008. CODEN ???? ISSN 0973-7766 (print), 2168-930X (electronic).

Kretschmer:2012:CPL

[KK12] Hildrun Kretschmer and Theo Kretschmer. Commemorating Professor Liu Zeyuan’s 50 years teaching: The decade from 1999– 2009 in collaboration with COLLNET. Collnet Journal of Scientometrics and Information Management, 6(2):185–214, 2012. CODEN ???? ISSN 0973-7766 (print), 2168-930X (electronic).

Kretschmer:2017:BCM

[KK17a] Hildrun Kretschmer and Theo Kretschmer. The beauty of crystals mirrored in collaboration networks of science. Collnet Journal of Scientometrics and Information Management, 11(1):73– 101, 2017. CODEN ???? ISSN 0973-7766 (print), 2168-930X

(electronic). URL http://www.tandfonline.com//doi/abs/
10.1080/09737766.2017.1303112.

Kumar:2017:PUL

[KK17b] D. Vinay Kumar and B. T. Sampath Kumar. Prevalence of URLs in Library and Information Science (LIS) literature: A citation analysis. Collnet Journal of Scientometrics and Information Management, 11(2):287–297, 2017. CODEN ???? ISSN 0973-7766 (print), 2168-930X (elec-

tronic). URL http://www.tandfonline.com//doi/abs/10.
1080/09737766.2017.1299390.

Kretschmer:2014:GSF

[KKS14] Hildrun Kretschmer, Theo Kretschmer, and Johannes Stegmann. Growth and structure formation of collaboration patterns obtained from the journals PNAS, Science and the Journal of Experimental Medicine. Collnet Journal of Scientometrics and Information Management, 8(1):1–15, 2014. CODEN ???? ISSN 0973-7766 (print), 2168-930X (electronic).

Kretschmer:2017:DEG

[KKYR17] Hildrun Kretschmer, Theo Kretschmer, Asundi A. Y., and I. K. Ravichandra Rao. Dr. Eugene Garfield: A humble homage to a great information scientist. Collnet Journal of Scientometrics and Information Management, 11(1):5–10, 2017. CODEN ???? ISSN 0973-7766 (print), 2168-930X

(electronic). URL http://www.tandfonline.com//doi/abs/
10.1080/09737766.2017.1344001.

Kretschmer:2017:RIC

[KLJ17] Hildrun Kretschmer, Jean-Charles Lamirel, and Pk Jain. Report of the 12th International Conference on Webometrics, Informetrics and Scientometrics (WIS) and 17th COLLNET Meeting, December 12–15, 2016, Nancy, France. Collnet Journal of Scientometrics and Information Management, 11(1):1– 4, 2017. CODEN ???? ISSN 0973-7766 (print), 2168-930X

(electronic). URL http://www.tandfonline.com//doi/abs/
10.1080/09737766.2017.1292665.

Khan:2020:ISW

[KMA20] Daud Khan, Basharat Ahmad Malik, and Naushad Ali. Identification of seminal works in the spectrum of knowledge management system through Reference Publication Year Spectroscopy (RPYS). Collnet Journal of Scientometrics and Information Management, 14(1):163–176, 2020. CODEN ???? ISSN 0973-7766 (print), 2168-930X (electronic). URL http:

//www.tandfonline.com/doi/abs/10.1080/09737766.2020.
1833691.

Kaur:2011:SAR

[KMG11] Har Kaur, Preeti Mahajan, and B. M. Gupta. Scientometric analysis of the research output: a study of Government Medical College & Hospital (GMCH), Chandigarh. Collnet Journal of Scientometrics and Information Management, 5(2):217–226, 2011. CODEN ???? ISSN 0973-7766 (print), 2168-930X (electronic).

Kretschmer:2011:CPE

[KN11] Hildrun Kretschmer and Kyungran Noh. Call for papers: Eighth International Conference on Webometrics, Informetrics and Scientometrics (WIS) & Thirteenth COLLNET Meeting. Collnet Journal of Scientometrics and Information Management, 5(2): 239–241, 2011. CODEN ???? ISSN 0973-7766 (print), 2168-930X (electronic).

Kokol:2019:FNF

[Kok19] Peter Kokol. Funded and non-funded research literature in software engineering in relation to country determinants. Collnet Journal of Scientometrics and Information Management, 13(1):

103–109, 2019. CODEN ???? ISSN 0973-7766 (print), 2168-930X

(electronic). URL http://www.tandfonline.com/doi/abs/10.
1080/09737766.2018.1560637.

Kretschmer:2019:ED

[Kre19] Hildrun Kretschmer. From Editor’s desk. Collnet Journal of Scientometrics and Information Management, 13(1):1– 2, 2019. CODEN ???? ISSN 0973-7766 (print), 2168-930X

(electronic). URL http://www.tandfonline.com/doi/abs/10.
1080/09737766.2019.1630895.

Khaparde:2010:EFL

[KS10] Vaishali Khaparde and Ramesh Sutar. Evaluation of the film library of films division in digital era: A case study. Collnet Journal of Scientometrics and Information Management, 4(2): 69–80, 2010. CODEN ???? ISSN 0973-7766 (print), 2168-930X (electronic).

Kumar:2019:INR

[KTCA19] Abhishek Kumar, Sanjay Tiwari, Ashish Kumar Chauhan, and Ramswaroop Ahirwar. Impact of NIRF on research publications: a study on top 20 (ranked) Indian universities. Collnet Journal of Scientometrics and Information Management, 13(2):219– 229, 2019. CODEN ???? ISSN 0973-7766 (print), 2168-930X

(electronic). URL http://www.tandfonline.com/doi/abs/10.
1080/09737766.2020.1741194.

Wan:2007:PIC

[kWhHR07] Jin kun Wan, Ping huan Hua, and Ronald Rousseau. The pure h-index: calculating an author’s h-index by taking co-authors into account. Collnet Journal of Scientometrics and Information Management, 1(2):1–5, 2007. CODEN ???? ISSN 0973-7766 (print), 2168-930X (electronic).

Kokol:2020:BRA

[KZZV20] Peter Kokol, Grega Zlahtic, Jernej Zavrsnik, and Helena Blazun Vosner. Buonocore research in adhesive dentistry: a remarkable sleeping paper. Collnet Journal of Scientometrics and Information Management, 14(2):211–217, 2020. CODEN ???? ISSN 0973-7766 (print), 2168-930X (electronic). URL http:

//www.tandfonline.com/doi/abs/10.1080/09737766.2021.
1906182.

Loan:2021:AAI

[LBN21] Fayaz Ahmad Loan, Bisma Bashir, and Nahida Nasreen. Applied Artificial Intelligence: a bibliometric study of an international journal. Collnet Journal of Scientometrics and Information Management, 15(1):27–45, 2021. CODEN ???? ISSN 0973-7766 (print), 2168-930X (electronic). URL http:

//www.tandfonline.com/doi/abs/10.1080/09737766.2021.
1938742.

Liwei:2015:SNA

[LC15] Zhang Liwei and Jiang Chunlin. Social network analysis and academic performance of the editorial board members for journals of library and information science. Collnet Journal of Scientometrics and Information Management, 9(2):131–143, 2015. CODEN ???? ISSN 0973-7766 (print), 2168-930X

(electronic). URL http://www.tandfonline.com/doi/abs/10.
1080/09737766.2015.1069947.

Levitt:2007:DAM

[Lev07] Jonathan M. Levitt. Developing and applying a metric for estimating trends in databases of articles. Collnet Journal of Scientometrics and Information Management, 1(1):17–21, 2007. CO-DEN ???? ISSN 0973-7766 (print), 2168-930X (electronic).

Lei:2018:GPB

[LFH18] Jing Lei, Hui-Zhen Fu, and Yuh-Shan Ho. A global perspective of bioaccumulation research using bibliometric analysis. Collnet Journal of Scientometrics and Information Management, 12(2): 327–341, 2018. CODEN ???? ISSN 0973-7766 (print), 2168-930X

(electronic). URL http://www.tandfonline.com/doi/abs/10.
1080/09737766.2018.1501923.

Librarian:2009:CCD

[LG09] Adarsh Bala Assistant Librarian and B. M. Gupta. S & T contribution of Chandigarh during 1998–2007: A case study. Collnet Journal of Scientometrics and Information Management, 3(2): 29–43, 2009. CODEN ???? ISSN 0973-7766 (print), 2168-930X (electronic).

Li:2014:BAN

[LGJ14] Jie Li, Xiaohong Guo, and Aleksandar Jovanovic. Bibliometrics analysis of nanosafety research. Collnet Journal of Scientometrics and Information Management, 8(2):437–455, 2014. CODEN ???? ISSN 0973-7766 (print), 2168-930X (electronic).

Lis:2009:CAL

[LLO09] Gholam Haidari Lis, Firoozeh Zareh Lis, and Farideh Osareh. Collaboration among Library and Information Science websites of Associations and Societies (LISAS), using webometric methods. Collnet Journal of Scientometrics and Information Management, 3(1):47–59, 2009. CODEN ???? ISSN 0973-7766 (print), 2168-930X (electronic).

Li:2016:ROJ

[LLZ16] Jie Li, Hui-Jie Li, and Lei Zhen. Research output of journal Technovation: A scientometric study. Collnet Journal of Scientometrics and Information Management, 10(2):255–272, 2016. CODEN ???? ISSN 0973-7766 (print), 2168-930X

(electronic). URL http://www.tandfonline.com/doi/abs/10.
1080/09737766.2016.1213968.

Maurya:2021:RPL

[LM21] Brady D. Lund and Sanjay Kumar Maurya. Research productivity of library and information science faculty in India and the United States: a comparison based on publications, citations and h-index. Collnet Journal of Scientometrics and Information Management, 15(1):89–105, 2021. CODEN ???? ISSN 0973-7766 (print), 2168-930X (elec-

tronic). URL http://www.tandfonline.com/doi/abs/10.
1080/09737766.2021.1936272.

Larsen:2008:SOI

[LMvI08] Peder Olesen Larsen, Isabelle Maye, and Markus von Ins. Scientific output and impact: Relative positions of China, Europe, India, Japan and the USA. Collnet Journal of Scientometrics and Information Management, 2(2):1–10, 2008. CODEN ???? ISSN 0973-7766 (print), 2168-930X (electronic).

Lateef:2016:GSC

[LOA16] A. Lateef, A. T. J. Ogunkunle, and G. O. Adigun. Google Scholar citation in retrospect: Visibility and contributions of African scholars. Collnet Journal of Scientometrics and Information Management, 10(2):219–236, 2016. CODEN ???? ISSN 0973-7766 (print), 2168-930X (electronic). URL http:

//www.tandfonline.com/doi/abs/10.1080/09737766.2016.
1213966.

Lathabai:2020:MIT

[LPGC20] Hiran H. Lathabai, Thara Prabhakaran, Susan George, and Manoj Changat. Mining interdisciplinary trajectories using multiple path analysis. Collnet Journal of Scientometrics and Information Management, 14(2):285–300, 2020. CODEN ???? ISSN 0973-7766 (print), 2168-930X (electronic). URL http:

//www.tandfonline.com/doi/abs/10.1080/09737766.2021.
1920066.

Lamirel:2014:AWC

[LR14] Jean-Charles Lamirel and David Reymond. Automatic websites classification and retrieval using websites communication signatures. Collnet Journal of Scientometrics and Information Management, 8(2):293–310, 2014. CODEN ???? ISSN 0973-7766 (print), 2168-930X (electronic).

Song:2019:CUP

[LS19] Chunjuan Luan and Bowen Song. Chinese university patenting, patent commercialization and economic growth: a provinciallevel analysis. Collnet Journal of Scientometrics and Information Management, 13(2):291–312, 2019. CODEN ???? ISSN 0973-7766 (print), 2168-930X (electronic). URL http:

//www.tandfonline.com/doi/abs/10.1080/09737766.2020.
1716644.

Markscheffel:2020:IAP

[Mar20] Bernd Markscheffel. Information about postponement of the 16^{th} International Conference on Webometrics, Informetrics and Scientometrics (WIS) and 21^{st} COLLNET Meeting. Collnet Journal of Scientometrics and Information Management, 14(1):1– 2, 2020. CODEN ???? ISSN 0973-7766 (print), 2168-930X

(electronic). URL http://www.tandfonline.com/doi/abs/10.
1080/09737766.2020.1818646.

Mayr:2008:EBE

[May08] Philipp Mayr. An evaluation of Bradfordizing effects. Collnet Journal of Scientometrics and Information Management, 2(2): 21–27, 2008. CODEN ???? ISSN 0973-7766 (print), 2168-930X (electronic).

Markscheffel:2012:TAS

[ME12] Bernd Markscheffel and Bastian Eine. The top 10 Alternative search Engines (ASE) — within selected categories ranked by Webometric indicators. Collnet Journal of Scientometrics and Information Management, 6(1):71–85, 2012. CODEN ???? ISSN 0973-7766 (print), 2168-930X (electronic).

Moazami:2015:NPA

[MES15] Afshin Moazami, Ashkan Ebadi, and Andrea Schiffauerova. A network perspective of academia–industry nanotechnology collaboration: A comparison of Canada and the United States. Collnet Journal of Scientometrics and Information Management, 9(2):263–293, 2015. CODEN ???? ISSN 0973-7766 (print),

2168-930X (electronic). URL http://www.tandfonline.com/
doi/abs/10.1080/09737766.2015.1069966.

Meyer:2007:BSM

[Mey07a] Martin Meyer. Biographical sketch of Martin Meyer. Collnet Journal of Scientometrics and Information Management, 1(1):v, 2007. CODEN ???? ISSN 0973-7766 (print), 2168-930X (electronic).

Meyer:2007:MP

[Mey07b] Martin Meyer. From Meyer’s pen. Collnet Journal of Scientometrics and Information Management, 1(1):vii, 2007. CODEN ???? ISSN 0973-7766 (print), 2168-930X (electronic).

Mo:2018:BHC

[MFH18] Ziwei Mo, Hui-Zhen Fu, and Yuh-Shan Ho. Bibliometrics of highly cited articles in the research field of volatile organic compounds. Collnet Journal of Scientometrics and Information Management, 12(2):309–326, 2018. CODEN ???? ISSN 0973-7766 (print), 2168-930X (electronic). URL http:

//www.tandfonline.com/doi/abs/10.1080/09737766.2018.
1524958.

Mohamadesmaeil:2012:PAK

[MG12] Sedigheh Mohamadesmaeil and Saeed Ghaffari. Promoting agriculture knowledge via public Web search engines: An experience by an Iranian librarian in response to agricultural queries. Collnet Journal of Scientometrics and Information Management, 6 (2):253–261, 2012. CODEN ???? ISSN 0973-7766 (print), 2168-930X (electronic).

Mallik:2018:SAR

[MG18] Ajoy Mallik and Biswajit Ghosh. Scientometric analysis of research advancement in graph theory and its applications. Collnet Journal of Scientometrics and Information Management, 12(2): 243–261, 2018. CODEN ???? ISSN 0973-7766 (print), 2168-930X

(electronic). URL http://www.tandfonline.com/doi/abs/10.
1080/09737766.2018.1453674.

Garg:2020:BGR

[MG20] Avijit Mahala and K. C. Garg. Bibliometrics of global research output in land degradation. Collnet Journal of Scientometrics and Information Management, 14(1):9–21, 2020. CODEN ???? ISSN 0973-7766 (print), 2168-930X (elec-

tronic). URL http://www.tandfonline.com/doi/abs/10.
1080/09737766.2020.1743789.

Motlagh:2017:KMC

[MH17] Fatemeh Helaliyan Motlagh and Mohammad Hassanzadeh. Knowledge management components in petrochemical industry: A case study. Collnet Journal of Scientometrics and Information Management, 11(1):11–23, 2017. CODEN ???? ISSN 0973-7766 (print), 2168-930X (electronic). URL http:

//www.tandfonline.com//doi/abs/10.1080/09737766.2016.
1239336.

Moghadami:2016:TYE

[MHS16] Maryam Moghadami, Mohammad Hassanzadeh, and Saeed Shokreian. Thirty years of Evidence-based Librarianship (EBL): Citation analysis and review. Collnet Journal of Scientometrics and Information Management, 10(2):209–218, 2016. CODEN ???? ISSN 0973-7766 (print), 2168-930X (elec-

tronic). URL http://www.tandfonline.com/doi/abs/10.
1080/09737766.2016.1213965.

Muthappan:2020:BAY

[MIS^{+}20] Sendhilkumar Muthappan, Kumaravel Ilangovan, Rathinavel Subramanian, Mantela Durairajan, and Rajalakshmi Elumalai. Bibliometric analysis of yoga scientific publications: an analysis of PubMed database 1948–2018. Collnet Journal of Scientometrics and Information Management, 14(2):301–309, 2020. CODEN ???? ISSN 0973-7766 (print), 2168-930X

(electronic). URL http://www.tandfonline.com/doi/abs/10.
1080/09737766.2021.1913976.

Mashroofa:2019:RTA

[MJC19] Mohamed Majeed Mashroofa, Mazuki Jusoh, and Karuthan Chinna. Research trend on the application of “e-learning adoption theory”: a scientometric study during 2000–2019, based on Web of Science and SCOPUS. Collnet Journal of Scientometrics and Information Management, 13(2):387–408, 2019. CODEN ???? ISSN 0973-7766 (print), 2168-930X

(electronic). URL http://www.tandfonline.com/doi/abs/10.
1080/09737766.2020.1729072.

Mohamadesmaeil:2012:WUE

[MK12] Sedigheh Mohamadesmaeil and Somaye Kazemi Koohbanani. Web usability evaluation of Iran National Library Website. Collnet Journal of Scientometrics and Information Management, 6 (1):161–174, 2012. CODEN ???? ISSN 0973-7766 (print), 2168-930X (electronic).

Markscheffel:2015:RIC

[MK15] Bernd Markscheffel and Hildrun Kretschmer. Report of The 10^{th} International Conference on Webometrics, Informetrics and Scientometrics (WIS). Collnet Journal of Scientometrics and Information Management, 9(1):1–4, 2015. CODEN ???? ISSN 0973-7766 (print), 2168-930X (electronic).

Markscheffel:2019:RIC

[MKP19] Bernd Markscheffel, Hildrun Kretschmer, and Pit Pichappan. Report of 14^{th} International Conference on Webometrics, Informetrics and Scientometrics (WIS) & 19th COLLNET Meeting 05 to 08 December 2018, University of Macau, Macau. Collnet Journal of Scientometrics and Information Management, 13(1): 3–6, 2019. CODEN ???? ISSN 0973-7766 (print), 2168-930X

(electronic). URL http://www.tandfonline.com/doi/abs/10.
1080/09737766.2019.1617229.

Markusova:2012:ICF

[MLA12] V. A. Markusova, A. N. Libkind, and E. Aversa. Impact of competitive funding on research output in Russia. Collnet Journal of Scientometrics and Information Management, 6(1):61–69, 2012. CODEN ???? ISSN 0973-7766 (print), 2168-930X (electronic).

Markusova:2014:BPT

[MLJM14] V. A. Markusova, A. N. Libkind, M. Jansz, and L. E. Mindeli. Bibliometric performance in two main research domains: The Russian Academy of Sciences and the higher education sector. Collnet Journal of Scientometrics and Information Management, 8(1):49–60, 2014. CODEN ???? ISSN 0973-7766 (print), 2168-930X (electronic).

Markusova:2013:RPF

[MLMJ13] V. A. Markusova, A. N. Libkind, L. E. Mindeli, and M. Jansz. Research performance by federal and national research universities and impact of competitive funding on their publication activity. Collnet Journal of Scientometrics and Information Management, 7(2):217–229, 2013. CODEN ???? ISSN 0973-7766 (print), 2168-930X (electronic).

Member:2009:IFI

[MMM09] Hamzehali Nourmohammadi Faculty Member, Abdolreza Noroozichakoli Faculty Member, and Mohammad Hassanzadeh Faculty Member. Impact factor in Institute for Scientific Information (ISI): Quality and quantity of scientific publications. Collnet Journal of Scientometrics and Information Management, 3(2):45–51, 2009. CODEN ???? ISSN 0973-7766 (print), 2168-930X (electronic).

Munnolli:2017:SSI

[MP17] Satish S. Munnolli and Shamprasad M. Pujar. Scientometric study of Indian cancer research based on Scopus database. Collnet Journal of Scientometrics and Information Management, 11 (2):201–214, 2017. CODEN ???? ISSN 0973-7766 (print), 2168-

930X (electronic). URL http://www.tandfonline.com//doi/
abs/10.1080/09737766.2017.1284729.

Mota:2017:MDC

[MPPH17] Fabio Batista Mota, Cláudio Damasceno Pinto, Julia Paranhos, and Lia Hasenclever. Mapping the ‘dynamic capabilities’ scientific landscape, 1990–2015: A bibliometric analysis. Collnet Journal of Scientometrics and Information Management, 11(2): 309–324, 2017. CODEN ???? ISSN 0973-7766 (print), 2168-

930X (electronic). URL http://www.tandfonline.com//doi/
abs/10.1080/09737766.2017.1306181.

Mahbuba:2016:NDA

[MR16a] Dilruba Mahbuba and Ronald Rousseau. New definitions and applications of year-based h-indices. Collnet Journal of Scientometrics and Information Management, 10(2):321–332, 2016. CODEN ???? ISSN 0973-7766 (print), 2168-930X

(electronic). URL http://www.tandfonline.com/doi/abs/10.
1080/09737766.2016.1229909.

Muthumari:2016:BAD

[MR16b] S. Muthumari and S. Raja. Bibliometric analysis of Defence Science Journal during 2005–2014: A study based on Scopus database. Collnet Journal of Scientometrics and Information Management, 10(2):273–287, 2016. CODEN ???? ISSN 0973-7766 (print), 2168-930X (electronic). URL http:

//www.tandfonline.com/doi/abs/10.1080/09737766.2016.
1213969.

Mondal:2018:FAC

[MR18] Dhiman Mondal and Nitai Raychoudhury. Foreign authorship collaboration in Indian LIS journals. Collnet Journal of Scientometrics and Information Management, 12(1):1–14, 2018. CODEN ???? ISSN 0973-7766 (print), 2168-930X

(electronic). URL http://www.tandfonline.com//doi/abs/
10.1080/09737766.2017.1332608.

Raychoudhury:2019:CIA

[MR19] Dhiman Mondal and Nitai Raychoudhury. Contribution of Indian authors in Scientometrics — an international journal during 1990–2017. Collnet Journal of Scientometrics and Information Management, 13(2):231–245, 2019. CODEN ???? ISSN 0973-7766 (print), 2168-930X (electronic). URL http:

//www.tandfonline.com/doi/abs/10.1080/09737766.2020.
1716642.

Md:2020:GQB

[MRAA20] Khadeeja M. N. Ansari Md, Safiqur Rahaman, and Hashem Hussein Al-Attas. Gauging the quality of behavioral science journals by using bibliometric indicators. Collnet Journal of Scientometrics and Information Management, 14(1):135–152, 2020. CODEN ???? ISSN 0973-7766 (print), 2168-930X

(electronic). URL http://www.tandfonline.com/doi/abs/10.
1080/09737766.2020.1819173.

Mosallaie:2021:DEA

[MRSE21] Shahab Mosallaie, Mahdi Rad, Andrea Schiffauerova, and Ashkan Ebadi. Discovering the evolution of artificial intelligence in cancer research using dynamic topic modeling. Collnet Journal of Scientometrics and Information Management, 15(2):225– 240, 2021. CODEN ???? ISSN 0973-7766 (print), 2168-930X

(electronic). URL http://www.tandfonline.com/doi/abs/10.
1080/09737766.2021.1958659.

Marshakova-Shaikevich:2007:MML

[MS07a] Irina Marshakova-Shaikevich. Memory and memories in lexical environment: bibliometric analysis of SSCI DB. Collnet Journal of Scientometrics and Information Management, 1(2):41–52, 2007. CODEN ???? ISSN 0973-7766 (print), 2168-930X (electronic).

Marshakova-Shaikevich:2007:VSC

[MS07b] Irina Marshakova-Shaikevich. The visualization of scientific collaboration of 15 ‘old’ and 10 ‘new’ EU countries in the field of social sciences. Collnet Journal of Scientometrics and Information Management, 1(1):9–16, 2007. CODEN ???? ISSN 0973-7766 (print), 2168-930X (electronic).

Mahawar:2012:BBP

[MS12a] K. L. Mahawar and O. P. Saini. Biblio-blogosphere: A platform for communication among LIS professionals. Collnet Journal of Scientometrics and Information Management, 6(2):301– 313, 2012. CODEN ???? ISSN 0973-7766 (print), 2168-930X (electronic).

Marshakova-Shaikevich:2012:PMJ

[MS12b] Irina Marshakova-Shaikevich. The Polish mathematical journals in world corpus of scientific journals presented in DBs JCR: bibliometric analysis. Collnet Journal of Scientometrics and Information Management, 6(2):279–299, 2012. CODEN ???? ISSN 0973-7766 (print), 2168-930X (electronic).

Marshakova-Shaikevich:2013:WNC

[MS13] Irina Marshakova-Shaikevich. World and national corpora of mathematical journals: case-study of Russia and Poland. Collnet Journal of Scientometrics and Information Management, 7 (2):277–291, 2013. CODEN ???? ISSN 0973-7766 (print), 2168-930X (electronic).

Meera:2014:ROU

[MS14] Meera and Surendra Kumar Sahu. Research output of University College of Medical Science, University of Delhi: A bibliometric study. Collnet Journal of Scientometrics and Information Management, 8(2):401–418, 2014. CODEN ???? ISSN 0973-7766 (print), 2168-930X (electronic).

Schrxfter:2021:CTS

[MS21] Bernd Markscheffel and Felix Schröter. Comparison of two science mapping tools based on software technical evaluation and bibliometric case studies. Collnet Journal of Scientometrics and Information Management, 15(2):365–396, 2021. CODEN ???? ISSN 0973-7766 (print), 2168-930X (elec-

tronic). URL http://www.tandfonline.com/doi/abs/10.
1080/09737766.2021.1960220.

Marshakova-Shaikevich:2008:JRW

[MSH08] Irina Marshakova-Shaikevich and Michael Heinz. Journal ranking of world scientific corpus by normalized impact factor. Collnet Journal of Scientometrics and Information Management, 2(2): 37–45, 2008. CODEN ???? ISSN 0973-7766 (print), 2168-930X (electronic).

Markscheffel:2008:RAS

[MT08] Bernd Markscheffel and Hendrik Thomas. Ranking approach for search engine result sets for a semi-automated generation of occurrences in topic maps. Collnet Journal of Scientometrics and Information Management, 2(1):91–98, 2008. CODEN ???? ISSN 0973-7766 (print), 2168-930X (electronic).

Markscheffel:2009:CCG

[MT09] Bernd Markscheffel and Hendrik Thomas. Coldlib — a concept for a global knowledge space for the Collnet community. Collnet Journal of Scientometrics and Information Management, 3(1): 11–16, 2009. CODEN ???? ISSN 0973-7766 (print), 2168-930X (electronic).

Mukherjee:2011:EIR

[Muk11] Bhaskar Mukherjee. Evaluation of Indian research performance in the emerging fields of environmental sciences. Collnet Journal of Scientometrics and Information Management, 5(1):81–98, 2011. CODEN ???? ISSN 0973-7766 (print), 2168-930X (electronic).

Must:2007:ICE

[Mus07] Ülle Must. International collaboration in EU 6^{th} Framework Programme: the case of India and China. Collnet Journal of Scientometrics and Information Management, 1(2):15–19, 2007. CODEN ???? ISSN 0973-7766 (print), 2168-930X (electronic).

Must:2011:IHR

[Mus11] Ülle Must. Implications on history research — does cooperation matter? Collnet Journal of Scientometrics and Information Management, 5(1):53–60, 2011. CODEN ???? ISSN 0973-7766 (print), 2168-930X (electronic).

Must:2014:IMA

[Mus14] Ülle Must. The impact of multi-authored papers: The case of a small country. Collnet Journal of Scientometrics and Information Management, 8(1):41–47, 2014. CODEN ???? ISSN 0973-7766 (print), 2168-930X (electronic).

Must:2017:BSS

[Mus17] Ülle Must. Bibliometric study on the sustainability of EU FP projects: Methodological aspects. Collnet Journal of Scientometrics and Information Management, 11(2):215–222, 2017. CODEN ???? ISSN 0973-7766 (print), 2168-930X (elec-

tronic). URL http://www.tandfonline.com//doi/abs/10.
1080/09737766.2017.1287822.

Mahesh:2012:WSB

[MW12] G. Mahesh and N. K. Wadhwa. Web of Science based ranking of Indian library and information science journals. Collnet Journal of Scientometrics and Information Management, 6(2):263–272, 2012. CODEN ???? ISSN 0973-7766 (print), 2168-930X (electronic).

Namdarian:2017:EST

[Nam17] Leila Namdarian. Evaluation of science, technology, and innovation (STI) in Iran. Collnet Journal of Scientometrics and Information Management, 11(2):253–271, 2017. CODEN ???? ISSN 0973-7766 (print), 2168-930X (electronic). URL http:

//www.tandfonline.com//doi/abs/10.1080/09737766.2017.
1321728.

Nazarovets:2018:HCP

[Naz18] Serhii Nazarovets. Highly cited papers of Ukrainian scientists written in collaboration: A bibliometric analysis (2011–2015). Collnet Journal of Scientometrics and Information Management, 12(1):35–47, 2018. CODEN ???? ISSN 0973-7766 (print), 2168-

930X (electronic). URL http://www.tandfonline.com//doi/
abs/10.1080/09737766.2017.1332613.

Negahban:2021:NAS

[Neg21] Mohammad Bagher Negahban. Network analysis and scientific mapping of the field of e-learning in Iran. Collnet Journal of Scientometrics and Information Management, 15(1):1– 8, 2021. CODEN ???? ISSN 0973-7766 (print), 2168-930X

(electronic). URL http://www.tandfonline.com/doi/abs/10.
1080/09737766.2020.1853489.

Nourmohammadi:2014:ASV

[NK14] Hamzehali Nourmohammadi and A. Keramatfar. Assessment of scientific visibility of Estonia on web: a study. Collnet Journal of Scientometrics and Information Management, 8(2):257–261, 2014. CODEN ???? ISSN 0973-7766 (print), 2168-930X (electronic).

Kumbar:2020:WCA

[NK20] B. Niveditha and Mallinath Kumbar. Web citation analysis of Library and Information Science and Communication and Media Studies journals: a comparative study. Collnet Journal of Scientometrics and Information Management, 14(2):335– 348, 2020. CODEN ???? ISSN 0973-7766 (print), 2168-930X

(electronic). URL http://www.tandfonline.com/doi/abs/10.
1080/09737766.2021.1915721.

Nourmohammadi:2015:RBW

[NKK15] Hamzehali Nourmohammadi, Mahdi Keramatfar, and Abdalsamad Keramatfar. The relation between Wikipedia’s size and scientific publication in a country. Collnet Journal of Scientometrics and Information Management, 9(1):83–92, 2015. CODEN ???? ISSN 0973-7766 (print), 2168-930X (electronic).

Nongrang:2016:BSB

[NL16] Khrawbok Nongrang and Bikika Laloo. Bibliometric study of biochemistry literature in North Eastern Hill University during 2000 to 2010. Collnet Journal of Scientometrics and Information Management, 10(2):197–207, 2016. CODEN ???? ISSN 0973-7766 (print), 2168-930X (electronic). URL http:

//www.tandfonline.com/doi/abs/10.1080/09737766.2016.
1213964.

Okagbue:2021:DCC

[OAdS21] Hilary I. Okagbue, Elvir M. Akhmetshin, and Jaime A. Teixeira da Silva. Distinct clusters of CiteScore and percentiles in top 1000 journals in Scopus. Collnet Journal of Scientometrics and Information Management, 15(1):133–143, 2021. CODEN ???? ISSN 0973-7766 (print), 2168-930X (elec-

tronic). URL http://www.tandfonline.com/doi/abs/10.
1080/09737766.2021.1934604.

Osareh:2014:BSG

[OAR14] Farideh Osareh, Elham Ahmadi, and Aref Riahi. The barriers and solutions for group work and scientific cooperation in Iran. Collnet Journal of Scientometrics and Information Management, 8(2):227–235, 2014. CODEN ???? ISSN 0973-7766 (print), 2168-930X (electronic).

Obermeier:2010:IPU

[OB10] Uwe Obermeier and Hannes Brauckmann. Interdisciplinary patterns of a university: Investigating collaboration using copublication network analysis. Collnet Journal of Scientometrics and Information Management, 4(1):29–40, 2010. CODEN ???? ISSN 0973-7766 (print), 2168-930X (electronic).

Osareh:2017:CSM

[OCK17] Farideh Osareh, Zohreh Cheraghi, and Atousa Koochak. Characterization of science mapping of Iranian cancer researchers’ co-authorship network in Web of Science service during 2000– 2013 time period. Collnet Journal of Scientometrics and Information Management, 11(2):223–234, 2017. CODEN ???? ISSN 0973-7766 (print), 2168-930X (electronic). URL http:

//www.tandfonline.com//doi/abs/10.1080/09737766.2017.
1292666.

Osareh:2014:CAN

[OKRS14] Farideh Osareh, Rouhallah Khademi, Mansoor Koohi Rostami, and Mansoureh Serati Shirazi. Co-authorship network structure analysis of Iranian researchers’ scientific outputs from 1991 to 2013 based on the Social Science Citation Index (SSCI). Collnet Journal of Scientometrics and Information Management, 8(2): 263–271, 2014. CODEN ???? ISSN 0973-7766 (print), 2168-930X (electronic).

Osareh:2008:VSI

[OM08] Farideh Osareh and Katherine W. McCain. Visualizing the structure of Iranian chemistry research in SciSearch using author cocitation technique. Collnet Journal of Scientometrics and Information Management, 2(1):37–43, 2008. CODEN ???? ISSN 0973-7766 (print), 2168-930X (electronic).

Osareh:2011:LLA

[OM11] Farideh Osareh and Esmaeel Mostafavi. Lotka’s Law and authorship distribution in computer science using Web of Science (WoS) during 1986–2009. Collnet Journal of Scientometrics and Information Management, 5(2):171–183, 2011. CODEN ???? ISSN 0973-7766 (print), 2168-930X (electronic).

Olechnicka:2010:MRS

[OP10] Agnieszka Olechnicka and Adam Ploszaj. Mapping the regional science performance. evidence from Poland. Collnet Journal of Scientometrics and Information Management, 4(1):21–27, 2010. CODEN ???? ISSN 0973-7766 (print), 2168-930X (electronic).

Ozel:2012:EKP

[ÖP12] Bülent Özel and Han Woo Park. Examining Korean political figures using co-word analysis in agreement with facial expressions in posted self-images. Collnet Journal of Scientometrics and Information Management, 6(1):43–60, 2012. CODEN ???? ISSN 0973-7766 (print), 2168-930X (electronic).

Osareh:2008:BSP

[Osa08] Farideh Osareh. Biographical sketch of Prof. Farideh Osareh. Collnet Journal of Scientometrics and Information Management, 2(2):iii, 2008. CODEN ???? ISSN 0973-7766 (print), 2168-930X (electronic).

Ozel:2011:BSD

[Öze11] Bülent Özel. Biographical sketch of Dr. Bülent Özel. Collnet Journal of Scientometrics and Information Management, 5(1): iii, 2011. CODEN ???? ISSN 0973-7766 (print), 2168-930X (electronic).

Pal:2016:ACC

[Pal16] Jiban K. Pal. Administering a cryptology centre by means of scientometric indicators. Collnet Journal of Scientometrics and Information Management, 10(1):97–123, 2016. CODEN ???? ISSN 0973-7766 (print), 2168-930X (electronic).

Pudovkin:2018:ACD

[PB18] Alexander I. Pudovkin and Lutz Bornmann. Approximation of citation distributions to the Poisson distribution. Collnet Journal of Scientometrics and Information Management, 12(1): 49–53, 2018. CODEN ???? ISSN 0973-7766 (print), 2168-

930X (electronic). URL http://www.tandfonline.com//doi/
abs/10.1080/09737766.2017.1332605.

Pearson:2020:PCW

[Pea20] William S. Pearson. Predictors of citations in written feedback on ESL writing research. Collnet Journal of Scientometrics and Information Management, 14(1):55–73, 2020. CODEN ???? ISSN 0973-7766 (print), 2168-930X (elec-

tronic). URL http://www.tandfonline.com/doi/abs/10.
1080/09737766.2020.1776651.

Pudovkin:2009:PRA

[PG09] A. I. Pudovkin and E. Garfield. Percentile rank and author superiority indexes for evaluating individual journal articles and the author’s overall citation performance. Collnet Journal of Scientometrics and Information Management, 3(2):3–10, 2009. CODEN ???? ISSN 0973-7766 (print), 2168-930X (electronic).

Parvathamma:2011:SPI

[PG11] N. Parvathamma and Devendra S. Gobbur. Scientific productivity and impact of Indian polymer scientists (1992–2006): An informetric study. Collnet Journal of Scientometrics and Information Management, 5(2):205–215, 2011. CODEN ???? ISSN 0973-7766 (print), 2168-930X (electronic).

Prithviraj:2015:WCT

[PK15] K. R. Prithviraj and B. T. Sampath Kumar. Web citation trends in Indian LIS journals: A citation analysis. Collnet Journal of Scientometrics and Information Management, 9(2):295– 310, 2015. CODEN ???? ISSN 0973-7766 (print), 2168-930X

(electronic). URL http://www.tandfonline.com/doi/abs/10.
1080/09737766.2015.1069968.

Pang:2009:DCA

[PKL09] Jie Pang, Hildrun Kretschmer, and Zeyuan Liu. Distribution of co-author pairs’ frequencies in energy science and technology. Collnet Journal of Scientometrics and Information Management, 3(2):11–18, 2009. CODEN ???? ISSN 0973-7766 (print), 2168-930X (electronic).

Pournaghi:2015:MRS

[PNA15] Roya Pournaghi and Leila Nemati-Anaraki. The mutual role of scientometrics and foresight — a review. Collnet Journal of Scientometrics and Information Management, 9(2):145– 160, 2015. CODEN ???? ISSN 0973-7766 (print), 2168-930X

(electronic). URL http://www.tandfonline.com/doi/abs/10.
1080/09737766.2015.1069950.

Papavlasopoulos:2009:TCB

[PPB09] Sozon Papavlasopoulos, Marios Poulos, and George Bokos. Towards constructing a bibliometric cited distance factor for an article using an Elman neural network. Collnet Journal of Scientometrics and Information Management, 3(1):17–29, 2009. CO-DEN ???? ISSN 0973-7766 (print), 2168-930X (electronic).

Pou:2021:AMT

[PRP21] K. R. Jolvis Pou, Vijaya Raghavan, and Muthukumaran Packirisamy. Applications of microfluidic technology in food sector: a bibliometric analysis. Collnet Journal of Scientometrics and Information Management, 15(2):259–285, 2021. CODEN ???? ISSN 0973-7766 (print), 2168-930X (elec-

tronic). URL http://www.tandfonline.com/doi/abs/10.
1080/09737766.2021.1989989.

Rahimmi:2020:DIF

[Rah20] Arman Rahimmi. Drawbacks to h-index, as a factor assessing the scientific impact and the scientific credit of a researcher. Collnet Journal of Scientometrics and Information Management, 14(2): 331–333, 2020. CODEN ???? ISSN 0973-7766 (print), 2168-930X

(electronic). URL http://www.tandfonline.com/doi/abs/10.
1080/09737766.2021.1906183.

Ram:2014:RPP

[Ram14] Shri Ram. Research profile of “polycystic ovary syndrome”: A bibliometric perspective of literatures from 1984–2013. Collnet Journal of Scientometrics and Information Management, 8(2): 311–327, 2014. CODEN ???? ISSN 0973-7766 (print), 2168-930X (electronic).

Rao:2012:RAN

[Rao12] I. K. Ravichandra Rao. Relations among the number of citations, references and authors: Revisited. Collnet Journal of Scientometrics and Information Management, 6(1):17–30, 2012. CO-DEN ???? ISSN 0973-7766 (print), 2168-930X (electronic).

Rao:2013:ED

[Rao13] I. K. Ravichandra Rao. From Editor’s desk. Collnet Journal of Scientometrics and Information Management, 7(1):i, 2013. CO-DEN ???? ISSN 0973-7766 (print), 2168-930X (electronic).

Rahimi:2013:CIO

[RC13] Shahin Rahimi and V. Chandrakumar. Citation impact of open access medical journals on scholarly communication: Web citation analysis. Collnet Journal of Scientometrics and Information Management, 7(2):247–259, 2013. CODEN ???? ISSN 0973-7766 (print), 2168-930X (electronic).

Rao:2011:EDa

[RK11a] I. K. Ravichandra Rao and Hildrun Kretschmer. From Editor’s desk. Collnet Journal of Scientometrics and Information Management, 5(1):i, 2011. CODEN ???? ISSN 0973-7766 (print), 2168-930X (electronic).

Rao:2011:EDb

[RK11b] I. K. Ravichandra Rao and Hildrun Kretschmer. From Editor’s desk. Collnet Journal of Scientometrics and Information Management, 5(2):i, 2011. CODEN ???? ISSN 0973-7766 (print), 2168-930X (electronic).

Rao:2012:EDb

[RK12] I. K. Ravichandra Rao and Hildrun Kretschmer. From Editor’s desk. Collnet Journal of Scientometrics and Information Management, 6(2):i, 2012. CODEN ???? ISSN 0973-7766 (print), 2168-930X (electronic).

Rao:2012:EDa

[RKÖG12] I. K. Ravichandra Rao, Hildrun Kretschmer, Bülent Özel, and Wolfgang Glänzel. From Editor’s desk. Collnet Journal of Scientometrics and Information Management, 6(1):i, 2012. CODEN ???? ISSN 0973-7766 (print), 2168-930X (electronic).

Rousseau:2008:RRD

[RL08] Ronald Rousseau and K. U. Leuven. Reflections on recent developments of the h-index and h-type indices. Collnet Journal of Scientometrics and Information Management, 2(1):1–8, 2008. CODEN ???? ISSN 0973-7766 (print), 2168-930X (electronic).

Rafols:2009:DDT

[RM09] Ismael Rafols and Martin Meyer. Disciplinary diversity and topic coherence: The case of hybrid nanomaterials research. Collnet Journal of Scientometrics and Information Management, 3(1): 79–88, 2009. CODEN ???? ISSN 0973-7766 (print), 2168-930X (electronic).

Ramkumar:2017:CNR

[RN17] S. Ramkumar and N. Narayanasamy. Collaboration and networking in research grant projects in speech, language and hearing sciences. Collnet Journal of Scientometrics and Information Management, 11(1):45–58, 2017. CODEN ???? ISSN 0973-7766 (print), 2168-930X (electronic). URL http:

//www.tandfonline.com//doi/abs/10.1080/09737766.2016.
1251558.

Rons:2011:IRC

[Ron11] Nadine Rons. Interdisciplinary research collaborations: Evaluation of a funding program. Collnet Journal of Scientometrics and Information Management, 5(1):17–32, 2011. CODEN ???? ISSN 0973-7766 (print), 2168-930X (electronic).

Rousseau:2012:FIS

[Rou12] Ronald Rousseau. Festschrifts in the information sciences, with special attention to Eugene Garfield’s festschrift “The Web of Knowledge”. Collnet Journal of Scientometrics and Information Management, 6(1):7–16, 2012. CODEN ???? ISSN 0973-7766 (print), 2168-930X (electronic).

Rousseau:2021:NNM

[Rou21] Ronald Rousseau. Naukometriya, Nalimov and Mul’chenko. Collnet Journal of Scientometrics and Information Management, 15 (1):213–224, 2021. CODEN ???? ISSN 0973-7766 (print), 2168-

930X (electronic). URL http://www.tandfonline.com/doi/
abs/10.1080/09737766.2021.1943042.

Putera:2021:CPP

[RP21] Tiurma Melissa Rakhel and Prakoso Bhairawa Putera. Corruption in public procurement: a bibliometric analysis. Collnet Journal of Scientometrics and Information Management, 15(2):397– 412, 2021. CODEN ???? ISSN 0973-7766 (print), 2168-930X

(electronic). URL http://www.tandfonline.com/doi/abs/10.
1080/09737766.2021.1989990.

Rousseau:2014:COI

[RR14] Brendan Rousseau and Ronald Rousseau. Calculating the outgrow index and similar structural indicators: A simple software program for visualizing outcomes. Collnet Journal of Scientometrics and Information Management, 8(1):31–40, 2014. CODEN ???? ISSN 0973-7766 (print), 2168-930X (electronic).

Raghavan:2015:MER

[RR15a] K. S. Raghavan and I. K. Ravichandra Rao. Mapping engineering research in India. Collnet Journal of Scientometrics and Information Management, 9(1):73–81, 2015. CODEN ???? ISSN 0973-7766 (print), 2168-930X (electronic).

Rao:2015:BJS

[RR15b] I. K. Ravichandra Rao and K. S. Raghavan. Collnet Journal of Scientometrics and Information Management (2007–2013): An analysis. Collnet Journal of Scientometrics and Information Management, 9(1):15–23, 2015. CODEN ???? ISSN 0973-7766 (print), 2168-930X (electronic).

Riahi:2014:ISI

[RRA14] Aref Riahi, Mohammad Amin Ghaneei Rod, and Elham Ahmadi. Iran’s scientific interactions and communications with the G8 countries. Collnet Journal of Scientometrics and Information Management, 8(2):217–225, 2014. CODEN ???? ISSN 0973-7766 (print), 2168-930X (electronic).

Rao:2017:QAS

[RRB17] I. K. Ravichandra Rao, K. S. Raghavan, and K. N. Bhargav. Quantitative analysis of scientific literature in forensic science. Collnet Journal of Scientometrics and Information Management, 11(1):59–72, 2017. CODEN ???? ISSN 0973-7766 (print), 2168-

930X (electronic). URL http://www.tandfonline.com//doi/
abs/10.1080/09737766.2016.1251561.

Rao:2008:DMA

[RS08] I. K. Ravichandra Rao and Bibhuti Bhusan Sahoo. Distributions of multiple authors: A case study of two journals (JASIST and Scientometrics). Collnet Journal of Scientometrics and Information Management, 2(1):27–36, 2008. CODEN ???? ISSN 0973-7766 (print), 2168-930X (electronic).

RayChoudhury:2011:MAR

[RS11] Nitai RayChoudhury and Juran Krishna Sarkhel. Mapping agricultural research in West Bengal — 1993–2007: A bibliometric study. Collnet Journal of Scientometrics and Information Management, 5(1):99–113, 2011. CODEN ???? ISSN 0973-7766 (print), 2168-930X (electronic).

Renjith:2018:HIG

[RS18] V. R. Renjith and I. Shihab. h-index of geology journals: A statistical analysis based on SCImago Journal and Country Ranking. Collnet Journal of Scientometrics and Information Management, 12(1):97–107, 2018. CODEN ???? ISSN 0973-7766 (print),

2168-930X (electronic). URL http://www.tandfonline.com//
doi/abs/10.1080/09737766.2017.1400753.

Riyanto:2019:DCM

[RSM^{+}19] Slamet Riyanto, Hendro Subagyo, Ekawati Marlina, Yaniasih Yaniasih, and Rodiah Rodiah. Development of collaboration matrix using co-occurrence and combinatoric system as scientometrics analysis. Collnet Journal of Scientometrics and Information Management, 13(2):247–267, 2019. CODEN ???? ISSN 0973-7766 (print), 2168-930X (electronic). URL http:

//www.tandfonline.com/doi/abs/10.1080/09737766.2020.
1723450.

Riahi:2015:IGI

[RSZ15] Aref Riahi, Fariba Sohbatiha, and Amin Zare. Investigation into growth of Iranian journals in Scopus database during 2000–2012. Collnet Journal of Scientometrics and Information Management, 9(1):37–46, 2015. CODEN ???? ISSN 0973-7766 (print), 2168-930X (electronic).

Ramakrishnan:2015:ICF

[RT15] J. Ramakrishnan and K. Thavamani. Indian contributions to the field of leptospirosis (2006–2013): A bibliometric study. Collnet Journal of Scientometrics and Information Management, 9(2): 235–249, 2015. CODEN ???? ISSN 0973-7766 (print), 2168-930X

(electronic). URL http://www.tandfonline.com/doi/abs/10.
1080/09737766.2015.1069962.

Roche:2014:HAR

[RVB^{+}14] Ivana Roche, Nathalie Vedovotto, Dominique Besagni, Claire François, Marianne Hörlesberger, and Edgar Schiebel. How to address the role of a journal as a knowledge transfer vector between disciplines? A case study relying on citation analysis. Collnet Journal of Scientometrics and Information Management, 8 (1):91–107, 2014. CODEN ???? ISSN 0973-7766 (print), 2168-930X (electronic).

Roche:2013:APA

[RVF^{+}13] Ivana Roche, Nathalie Vedovotto, Claire François, Dominique Besagni, Marianne Hörlesberger, and Edgar Schiebel. From applicateness to potential applicability: An informetric approach based on researchers’ scientific productions. Collnet Journal of Scientometrics and Information Management, 7(2):191–203, 2013. CODEN ???? ISSN 0973-7766 (print), 2168-930X (electronic).

Rousseau:2018:URM

[RY18] Ronald Rousseau and Sulan Yan. An under-recognized metatable: Nancy Clark’s table tamer. Collnet Journal of Scientometrics and Information Management, 12(2):233–241, 2018. CODEN ???? ISSN 0973-7766 (print), 2168-930X (elec-

tronic). URL http://www.tandfonline.com/doi/abs/10.
1080/09737766.2018.1453673.

Senthilkumaran:2007:CNA

[SA07] P. Senthilkumaran and A. Amudhavalli. Cross-national analysis on spices research in the Asian countries. Collnet Journal of Scientometrics and Information Management, 1(2):27–32, 2007. CODEN ???? ISSN 0973-7766 (print), 2168-930X (electronic).

Sangam:2016:GVS

[SA16] S. L. Sangam and Uma Arali. Growth versus scientific collaboration in the field of genetics: A scientometrics analysis. Collnet Journal of Scientometrics and Information Management, 10(1): 9–19, 2016. CODEN ???? ISSN 0973-7766 (print), 2168-930X (electronic).

Safavi:2014:BRP

[Saf14] Zainab Safavi. Barriers to research from the perspective of faculty members of knowledge and information science: a case study of public universities in Tehran. Collnet Journal of Scientometrics and Information Management, 8(2):281–291, 2014. CODEN ???? ISSN 0973-7766 (print), 2168-930X (electronic).

Sangam:2010:BSP

[San10] S. L. Sangam. Biographical sketch of Professor S. L. Sangam. Collnet Journal of Scientometrics and Information Management, 4(1):iii, 2010. CODEN ???? ISSN 0973-7766 (print), 2168-930X (electronic).

Sangam:2013:SAG

[SAPM13] S. L. Sangam, Uma Arali, C. G. Patil, and M. N. Megeri. Scientometrics analysis of genetics literature. Collnet Journal of Scientometrics and Information Management, 7(2):173–190, 2013. CODEN ???? ISSN 0973-7766 (print), 2168-930X (electronic).

Singh:2014:CCL

[SB14] K. P. Singh and Bebi. Characteristics and characterization of literature used by the researchers in geography: A citation analysis of Ph.D. theses. Collnet Journal of Scientometrics and Information Management, 8(2):341–351, 2014. CODEN ???? ISSN 0973-7766 (print), 2168-930X (electronic).

Stumpf:2012:SSB

[SCdM^{+}12] Ida Regina Chittó Stumpf, Sônia Elisa Caregnato, Ana Maria Mielniczuk de Moura, Samile Andréa de Souza Vanz, and Rosely de Andrade Vargas. Science in South Brazil: Production overview between 2000 and 2010. Collnet Journal of Scientometrics and Information Management, 6(1):119–130, 2012. CODEN ???? ISSN 0973-7766 (print), 2168-930X (electronic).

Scientist:2014:WRA

[Sci14] I. K. Ravichandra Rao Visiting Scientist. Weak relations among the impact factors, number of citations, references and authors. Collnet Journal of Scientometrics and Information Management, 8(1):17–30, 2014. CODEN ???? ISSN 0973-7766 (print), 2168-930X (electronic).

Savanur:2014:LLA

[SDK14] Kiran Savanur, S. Gayathri Devi, and P. V. Konnur. Lotka’s Law and authorship distribution in the journal of ‘Columbia Law Review’. Collnet Journal of Scientometrics and Information Management, 8(1):193–208, 2014. CODEN ???? ISSN 0973-7766 (print), 2168-930X (electronic).

Singh:2010:MIE

[SG10] Yogendra Singh and B. M. Gupta. Mapping of Indian engineering research using quantitative indicators, 1999–2008. Collnet Journal of Scientometrics and Information Management, 4(2): 1–12, 2010. CODEN ???? ISSN 0973-7766 (print), 2168-930X (electronic).

Shuang:2016:CSH

[SGP16] Yong-Qiang Shuang, Wei-Hao Gao, and Meng-Han Pang. A comparative study on the highly cited papers of the top eight engineering universities. Collnet Journal of Scientometrics and Information Management, 10(2):311–320, 2016. CODEN ???? ISSN 0973-7766 (print), 2168-930X (elec-

tronic). URL http://www.tandfonline.com/doi/abs/10.
1080/09737766.2016.1229908.

Shafi:2010:BSP

[Sha10] S. M. Shafi. Biographical sketch of Professor S. M. Shafi. Collnet Journal of Scientometrics and Information Management, 4(2): iv, 2010. CODEN ???? ISSN 0973-7766 (print), 2168-930X (electronic).

Shatberashvili:2019:MEB

[Sha19] Oleg Shatberashvili. Method of express bibliometric analysis of scientific diasporas. Collnet Journal of Scientometrics and Information Management, 13(1):139–145, 2019. CODEN ???? ISSN 0973-7766 (print), 2168-930X (electronic). URL http:

//www.tandfonline.com/doi/abs/10.1080/09737766.2019.
1568364.

Shukla:2019:IRO

[Shu19] Ravi Shukla. Indian research output on genetic disorder publication using the Scopus database: a scientometric study. Collnet Journal of Scientometrics and Information Management, 13(1): 91–102, 2019. CODEN ???? ISSN 0973-7766 (print), 2168-930X

(electronic). URL http://www.tandfonline.com/doi/abs/10.
1080/09737766.2018.1550044.

Singh:2015:LPW

[Sin15] Malkeet Singh. Literature published on Web 2.0 technologies in libraries: A bibliometric study. Collnet Journal of Scientometrics and Information Management, 9(2):251–262, 2015. CODEN ???? ISSN 0973-7766 (print), 2168-930X

(electronic). URL http://www.tandfonline.com/doi/abs/10.
1080/09737766.2015.1069963.

Sastry:2016:ARM

[SJM16] Chandramouli Shama Sastry, Darshan S. Jagaluru, and Kavi Mahesh. Author ranking in multi-author collaborative networks. Collnet Journal of Scientometrics and Information Management, 10(1):21–40, 2016. CODEN ???? ISSN 0973-7766 (print), 2168-930X (electronic).

Saha:2016:DEF

[SJMN16] Snehanshu Saha, Neelam Jangid, Archana Mathur, and Anand M. Narsimhamurthy. DSRS: Estimation and forecasting of journal influence in the science and technology domain via a lightweight quantitative approach. Collnet Journal of Scientometrics and Information Management, 10(1):41–70, 2016. CODEN ???? ISSN 0973-7766 (print), 2168-930X (electronic).

Sheshrao:2011:CAP

[SK11] Jadhav Vandana Sheshrao and V. S. Khaparde. Citation analysis of Ph.D. theses on physics submitted to Dr. Babasaheb Ambedkar Marathwada University. Collnet Journal of Scientometrics and Information Management, 5(1):115–127, 2011. CODEN ???? ISSN 0973-7766 (print), 2168-930X (electronic).

Sajana:2017:RCS

[SK17] C. Sajana and M. Krishnamurthy. Research contributions of space technology scientists: A scientometric study. Collnet Journal of Scientometrics and Information Management, 11(2): 325–339, 2017. CODEN ???? ISSN 0973-7766 (print), 2168-

930X (electronic). URL http://www.tandfonline.com//doi/
abs/10.1080/09737766.2017.1312788.

Sun:2008:IAR

[SKNN08] Yuan Sun, Sumio Kakinuma, Masamitsu Negishi, and Masaki Nisizawa. Internationalizing academic research activities in Japan. Collnet Journal of Scientometrics and Information Management, 2(2):11–19, 2008. CODEN ???? ISSN 0973-7766 (print), 2168-930X (electronic).

Sangam:2009:RCP

[SM09] S. L. Sangam and Meera. Research collaboration pattern in Indian contributions to chemical sciences. Collnet Journal of Scientometrics and Information Management, 3(1):39–45, 2009. CO-DEN ???? ISSN 0973-7766 (print), 2168-930X (electronic).

Sangam:2014:MIS

[SM14] S. L. Sangam and Shivaranjini S. Mogali. Mapping of Indian social science literature. Collnet Journal of Scientometrics and Information Management, 8(1):61–80, 2014. CODEN ???? ISSN 0973-7766 (print), 2168-930X (electronic).

Shukla:2018:RPS

[SM18] Akhandanand Shukla and Sanjay Kumar Maurya. Research performance of South Asian Association for Regional Cooperation (SAARC) countries in library and information science: A scientometric analysis. Collnet Journal of Scientometrics and Information Management, 12(1):73–81, 2018. CODEN ???? ISSN 0973-7766 (print), 2168-930X (elec-

tronic). URL http://www.tandfonline.com//doi/abs/10.
1080/09737766.2017.1354478.

Sangam:2015:SPG

[SMA15] S. L. Sangam, Devika Madalli, and Uma B. Arali. Scientometrics profile of global genetics literature as seen through PubMed. Collnet Journal of Scientometrics and Information Management, 9(2):175–192, 2015. CODEN ???? ISSN 0973-7766 (print),

2168-930X (electronic). URL http://www.tandfonline.com/
doi/abs/10.1080/09737766.2015.1069956.

Sangam:2008:GPI

[SMM08] S. L. Sangam, Meera, and M. N. Megeri. Growth pattern of Indian chemical science literature: A scientometric analysis. Collnet Journal of Scientometrics and Information Management, 2 (1):99–110, 2008. CODEN ???? ISSN 0973-7766 (print), 2168-930X (electronic).

Sun:2011:CER

[SNK11] Yuan Sun, Masaki Nisizawa, and Sumio Kakinuma. Competitive external research funding as an indicator in university research evaluation: Fractional counting of number of funding applications. Collnet Journal of Scientometrics and Information Management, 5(2):157–169, 2011. CODEN ???? ISSN 0973-7766 (print), 2168-930X (electronic).

Pandey:2020:MGP

[SP20] Sidhartha Sahoo and Shriram Pandey. Measuring growth pattern and publications trends of storm surges research during (1999–2019): a scientometric study. Collnet Journal of Scientometrics and Information Management, 14(2):237–256, 2020. CODEN ???? ISSN 0973-7766 (print), 2168-930X

(electronic). URL http://www.tandfonline.com/doi/abs/10.
1080/09737766.2021.1918031.

Syamili:2017:DAC

[SR17] C. Syamili and R. V. Rekha. Do altmetric correlate with citation?: A study based on PLOS ONE journal. Collnet Journal of Scientometrics and Information Management, 11(1):103– 117, 2017. CODEN ???? ISSN 0973-7766 (print), 2168-930X

(electronic). URL http://www.tandfonline.com//doi/abs/
10.1080/09737766.2016.1260815.

Satija:2007:LRV

[SS07] Mp Satija and Aya Singh. Are laws of Ranganathan valid in virtual libraries? Collnet Journal of Scientometrics and Information Management, 1(1):27–34, 2007. CODEN ???? ISSN 0973-7766 (print), 2168-930X (electronic).

Sangam:2010:EGS

[SS10] S. L. Sangam and K. Savanur. Eugene Garfield: A scientometric portrait. Collnet Journal of Scientometrics and Information Management, 4(1):41–51, 2010. CODEN ???? ISSN 0973-7766 (print), 2168-930X (electronic).

S:2014:INR

[SS14] Sri Amudha S. and R. Sevukan. Indian neuroscience research, 1999–2013: A scientometric analysis. Collnet Journal of Scientometrics and Information Management, 8(2):329–340, 2014. CODEN ???? ISSN 0973-7766 (print), 2168-930X (electronic).

Sangam:2019:CCG

[SS19] S. L. Sangam and K. S. Savitha. Climate change and global warming: a scientometric study. Collnet Journal of Scientometrics and Information Management, 13(1):199–212, 2019. CODEN ???? ISSN 0973-7766 (print), 2168-930X

(electronic). URL http://www.tandfonline.com/doi/abs/10.
1080/09737766.2019.1598001.

Stoupas:2021:WUR

[SSKM21] Georgios Stoupas, Antonis Sidiropoulos, Dimitrios Katsaros, and Yannis Manolopoulos. When universities rise (rank) high into the skyline. Collnet Journal of Scientometrics and Information Management, 15(2):241–258, 2021. CODEN ???? ISSN 0973-7766 (print), 2168-930X (electronic). URL http:

//www.tandfonline.com/doi/abs/10.1080/09737766.2021.
1955419.

Sajovic:2018:BSM

[STP18] Irena Sajovic, Helena Gabrijelcic Tomc, and Bojana Boh Podgornik. Bibliometric study and mapping of a journal in the field of visualization and computer graphics. Collnet Journal of Scientometrics and Information Management, 12(2):263– 287, 2018. CODEN ???? ISSN 0973-7766 (print), 2168-930X

(electronic). URL http://www.tandfonline.com/doi/abs/10.
1080/09737766.2018.1453677.

Sudhier:2010:BLS

[Sud10] K. G. Sudhier. Bradford’s law of scattering revisited: A study based on the references in doctoral theses in the area of physics. Collnet Journal of Scientometrics and Information Management, 4(2):35–47, 2010. CODEN ???? ISSN 0973-7766 (print), 2168-930X (electronic).

Sun:2007:CRN

[Sun07] Yuan Sun. Collaborative research networks of Japanese universities: Bibliometric trends. Collnet Journal of Scientometrics and Information Management, 1(2):7–13, 2007. CODEN ???? ISSN 0973-7766 (print), 2168-930X (electronic).

Suk:2019:HRC

[SWH19] Fat-Moon Suk, Ming-Huang Wang, and Yuh-Shan Ho. A historical review of classic Helicobacter pylori publications in Science Citation Index expanded. Collnet Journal of Scientometrics and Information Management, 13(2):313–330, 2019. CODEN ???? ISSN 0973-7766 (print), 2168-930X

(electronic). URL http://www.tandfonline.com/doi/abs/10.
1080/09737766.2020.1716645.

Tijssen:2008:WMT

[Tij08] Robert J. W. Tijssen. Are we moving towards an integrated European research area? Some macro-level bibliometric perspectives. Collnet Journal of Scientometrics and Information Management, 2(1):19–25, 2008. CODEN ???? ISSN 0973-7766 (print), 2168-930X (electronic).

Tripathi:2018:OAK

[TKSB18] Manorama Tripathi, Sunil Kumar, S. K. Sonker, and Parveen Babbar. Occurrence of author keywords and keywords plus in social sciences and humanities research: a preliminary study. Collnet Journal of Scientometrics and Information Management, 12(2):215–232, 2018. CODEN ???? ISSN 0973-7766 (print),

2168-930X (electronic). URL http://www.tandfonline.com/
doi/abs/10.1080/09737766.2018.1436951.

Usman:2019:BAN

[UE19] Muhammad Kabiru Usman and Ozioma Ewulum. A bibliometric analysis of Nigeria’s Library and Information Sciences literature: a study of Journal of Applied Information Science and Technology. Collnet Journal of Scientometrics and Information Management, 13(1):53–64, 2019. CODEN ???? ISSN 0973-7766 (print),

2168-930X (electronic). URL http://www.tandfonline.com/
doi/abs/10.1080/09737766.2018.1541042.

Usman:2017:UTT

[UY17] M. K. Usman and A. D. Yahaya. The usability traits of theses and dissertations as sources of information to agriculture students. Collnet Journal of Scientometrics and Information Management, 11(1):119–131, 2017. CODEN ???? ISSN 0973-7766 (print),

2168-930X (electronic). URL http://www.tandfonline.com//
doi/abs/10.1080/09737766.2016.1266144.

Vaughan:2008:NFI

[Vau08] Liwen Vaughan. A new frontier of informetric and Webometric research: Mining Web usage data. Collnet Journal of Scientometrics and Information Management, 2(2):29–35, 2008. CO-DEN ???? ISSN 0973-7766 (print), 2168-930X (electronic).

Vaughan:2012:ADS

[Vau12] Liwen Vaughan. An alternative data source for Web hyperlink analysis: “Sites linking in” at Alexa Internet. Collnet Journal of Scientometrics and Information Management, 6(1):31–42, 2012. CODEN ???? ISSN 0973-7766 (print), 2168-930X (electronic).

Verma:2017:RNI

[VB17] Manoj Kumar Verma and Krishna Brahma. Ranking of National Institutes of Technology (NITs) of northeast region of India on the basis of web impact factor. Collnet Journal of Scientometrics and Information Management, 11(2):235–248, 2017. CODEN ???? ISSN 0973-7766 (print), 2168-930X

(electronic). URL http://www.tandfonline.com//doi/abs/
10.1080/09737766.2017.1292667.

Verma:2009:BSI

[Ver09] R. K. Verma. Bibliometric study of Indian S & T publications as reflected in Indian science abstracts. Collnet Journal of Scientometrics and Information Management, 3(2):19–28, 2009. CO-DEN ???? ISSN 0973-7766 (print), 2168-930X (electronic).

Vutsova:2021:IAR

[VHA21] Albena Vutsova, Todor Hristov, and Martina Arabadzhieva. Impact assessment of research evaluation in Bulgaria. Collnet Journal of Scientometrics and Information Management, 15(2):341– 364, 2021. CODEN ???? ISSN 0973-7766 (print), 2168-930X

(electronic). URL http://www.tandfonline.com/doi/abs/10.
1080/09737766.2021.1962767.

Velmurugan:2015:SOA

[VR15] Chandran Velmurugan and Natarajan Radhakrishnan. Scientometric observations of authorship trends and collaborative research on DESIDOC Journal of Library and Information Technology. Collnet Journal of Scientometrics and Information Management, 9(2):193–204, 2015. CODEN ???? ISSN 0973-7766 (print),

2168-930X (electronic). URL http://www.tandfonline.com/
doi/abs/10.1080/09737766.2015.1069957.

Vijayabanu:2019:BES

[VT19] C. Vijayabanu and C. Therasa. Bibliometric evidence: Statistical outcome of personality — psychology research using Scopus database from 1860 to 2017. Collnet Journal of Scientometrics and Information Management, 13(1):183–197, 2019. CODEN ???? ISSN 0973-7766 (print), 2168-930X

(electronic). URL http://www.tandfonline.com/doi/abs/10.
1080/09737766.2019.1583837.

Walke:2007:MSR

[WDG07] Rajpal Walke, S. M. Dhawan, and B. M. Gupta. Materials science research in India (1993–2001): A scientometric study. Collnet Journal of Scientometrics and Information Management, 1(2): 33–40, 2007. CODEN ???? ISSN 0973-7766 (print), 2168-930X (electronic).

Wang:2019:BAD

[WH19] Xiaoru Wang and Yuh-Shan Ho. A bibliometric analysis of dance performance reviews in the dance category of the Web of Science. Collnet Journal of Scientometrics and Information Management, 13(1):79–90, 2019. CODEN ???? ISSN 0973-7766 (print), 2168-

930X (electronic). URL http://www.tandfonline.com/doi/
abs/10.1080/09737766.2018.1550036.

Wadhwa:2010:CAP

[WKK10] Narendar Kumar Wadhwa, Hildrun Kretschmer, and Theo Kretschmer. Co-author pairs’ frequencies of the national physical laboratory India. Collnet Journal of Scientometrics and Information Management, 4(2):49–55, 2010. CODEN ???? ISSN 0973-7766 (print), 2168-930X (electronic).

Wang:2020:BAK

[WWH20] Yi-Fen Wang, Ming-Huang Wang, and Yuh-Shan Ho. A bibliometric analysis of KSHV/HHV8 research. Collnet Journal of Scientometrics and Information Management, 14(2):219–235, 2020. CODEN ???? ISSN 0973-7766 (print), 2168-930X

(electronic). URL http://www.tandfonline.com/doi/abs/10.
1080/09737766.2021.1920065.

Wahid:2021:GLS

[WWT21] Nazia Wahid, Nosheen Fatima Warraich, and Muzammil Tahira. Group level scientometric analysis of Pakistani authors. Collnet Journal of Scientometrics and Information Management, 15(2): 287–304, 2021. CODEN ???? ISSN 0973-7766 (print), 2168-930X

(electronic). URL http://www.tandfonline.com/doi/abs/10.
1080/09737766.2021.1960219.

Wang:2014:EKN

[WX14] H. Wang and J. Xiao. Effects of knowledge network modularity on the utility of inventions in semiconductor industry. Collnet Journal of Scientometrics and Information Management, 8(1): 141–152, 2014. CODEN ???? ISSN 0973-7766 (print), 2168-930X (electronic).

Wani:2017:RES

[WZ17] Zahid Ashraf Wani and Tazeem Zainab. A review of eminence of scientometric indicators in scientific research productivity. Collnet Journal of Scientometrics and Information Management, 11 (2):273–285, 2017. CODEN ???? ISSN 0973-7766 (print), 2168-

930X (electronic). URL http://www.tandfonline.com//doi/
abs/10.1080/09737766.2017.1306179.

Ye:2017:HBI

[YBL17] Fred Y. Ye, Lutz Bornmann, and Loet Leydesdorff. h-based I3-type multivariate vectors: multidimensional indicators of publication and citation scores. Collnet Journal of Scientometrics and Information Management, 11(1):153–171, 2017. CODEN ???? ISSN 0973-7766 (print), 2168-930X (elec-

tronic). URL http://www.tandfonline.com//doi/abs/10.
1080/09737766.2017.1284954.

Yang:2013:CCI

[YLC13] Kiduk Yang, Jongwook Lee, and Seon-Heui Choi. Comparison of citation indexes in Korea: An exploratory study. Collnet Journal of Scientometrics and Information Management, 7(2):231– 245, 2013. CODEN ???? ISSN 0973-7766 (print), 2168-930X (electronic).

Yue:2020:RIC

[YLM20] Chen Yue, Jean-Charles Lamirel, and Bernd Markscheffel. Report of the 15^{th} International Conference on Webometrics, Informetrics and Scientometrics (WIS) and the 20^{th} COLL-NET Meeting. Collnet Journal of Scientometrics and Information Management, 14(1):203–209, 2020. CODEN ???? ISSN 0973-7766 (print), 2168-930X (electronic). URL http:

//www.tandfonline.com/doi/abs/10.1080/09737766.2020.
1838767.

Yang:2014:RAE

[YLX14] Z. K. Yang, D. M. Lin, and M. Z. Xu. The re-applicability explore of Lotka’s Law in patent documents. Collnet Journal of Scientometrics and Information Management, 8(1):183–191, 2014. CO-DEN ???? ISSN 0973-7766 (print), 2168-930X (electronic).

Yue:2015:IBA

[YLZS15] Chen Yue, Zhang Liwei, Wang Zhiqi, and Liu Shengbo. Influential bloggers and active bloggers on ScienceNet: An analysis of popular blogs. Collnet Journal of Scientometrics and Information Management, 9(2):121–130, 2015. CODEN ???? ISSN 0973-7766 (print), 2168-930X (electronic). URL http:

//www.tandfonline.com/doi/abs/10.1080/09737766.2015.
1069945.

Sun:2019:BMT

[YS19] Bao-Zhong Yuan and Jie Sun. Bibliometric and mapping of top papers in the subject category of green and sustainable science and technology based on ESI. Collnet Journal of Scientometrics and Information Management, 13(2):269–289, 2019. CODEN ???? ISSN 0973-7766 (print), 2168-930X

(electronic). URL http://www.tandfonline.com/doi/abs/10.
1080/09737766.2020.1716643.

Sun:2020:BAR

[YS20] Bao-Zhong Yuan and Jie Sun. Bibliometric analysis of research on the maize based on top papers during 2009–2019. Collnet Journal of Scientometrics and Information Management, 14(1): 75–92, 2020. CODEN ???? ISSN 0973-7766 (print), 2168-930X

(electronic). URL http://www.tandfonline.com/doi/abs/10.
1080/09737766.2020.1787110.

Zeyuan:2009:BSP

[Zey09] Liu Zeyuan. Biographical sketch of Prof. Liu Zeyuan. Collnet Journal of Scientometrics and Information Management, 3(1): iii, 2009. CODEN ???? ISSN 0973-7766 (print), 2168-930X (electronic).

Zavaraqi:2012:SSS

[ZF12] Rasoul Zavaraqi and Gholam-Reza Fadaie. Scientometrics or science of science: quantitative, qualitative or mixed one. Collnet Journal of Scientometrics and Information Management, 6(2):

273–278, 2012. CODEN ???? ISSN 0973-7766 (print), 2168-930X (electronic).

Zitt:2013:VCB

[Zit13] Michel Zitt. Variability of citation behavior between scientific fields and the normalization problem: The “citing-side” normalization in context. Collnet Journal of Scientometrics and Information Management, 7(1):55–67, 2013. CODEN ???? ISSN 0973-7766 (print), 2168-930X (electronic).

Zhang:2019:SCA

[ZKJK19] Xiao Juan Zhang, Farhan Khan, Xiang Jinpeng, and Kashif Ullah Khan. Study of cognitive and affected trust in knowledge sharing evidence from Chinese firms — a review paper. Collnet Journal of Scientometrics and Information Management, 13(1): 147–165, 2019. CODEN ???? ISSN 0973-7766 (print), 2168-930X

(electronic). URL http://www.tandfonline.com/doi/abs/10.
1080/09737766.2019.1569825.

Zlahtic:2019:DIS

[ZKZ19] Grega Zlahtic, Peter Kokol, and Bojan Zlahtic. Discrepancies in identifying sleeping papers in Scopus and Web of science: The case of “software engineering”. Collnet Journal of Scientometrics and Information Management, 13(2):339–344, 2019. CODEN ???? ISSN 0973-7766 (print), 2168-930X

(electronic). URL http://www.tandfonline.com/doi/abs/10.
1080/09737766.2020.1720549.

Zahedi:2012:ICP

[ZS12] Zohreh Zahedi and Mansoureh Serati Shirazi. An investigation of citations pattern to Iranian’s English journals using Google Scholar & ISI: Do they differ? Collnet Journal of Scientometrics and Information Management, 6(1):87–96, 2012. CODEN ???? ISSN 0973-7766 (print), 2168-930X (electronic).